### v1

# Dressed to Gamble: How Poker Drives the Dynamics of Wearables and Visits on Decentraland’s Social Virtual World

Abstract
Decentraland is a blockchain-based social virtual world touted to be a creative space owned by its community, unlike previous virtual worlds. Its users can create and publish wearables, virtual garments to customize avatars, which can be then sold or given away via the blockchain. Decentral Games (DG), a single project owning two prominent in-world casinos, has by far created the most wearables, with these being necessary to earn cryptocurrency in their flagship game ICE Poker. We thus present a comprehensive study that investigates how DG and ICE Poker influence the overall dynamics of Decentraland wearables and in-world visits. To this end, we analyzed 5.9 million wearable transfers made on the Polygon blockchain (and related sales) over a two-year period, and 677 million log events of in-world user positions in an overlapping 10-month period. We found that the influence of DG and Ice Poker is not only significant, but also substantial for transfers and sales monetary value of wearables, and very large for daily unique visitors and time spent in the virtual world. Despite several alternative in-world economic and artistic initiatives in Decentraland, some of which have attracted much attention from the general public, online poker appears to be the main driver of the analyzed dynamics. Our work thus contributes to the current understanding of user behavior in social virtual worlds and it is among the first to study the emerging phenomenon of blockchain-based online gambling in virtual spaces.
keywords:
social virtual worlds, Metaverse, online gambling, NFT, blockchain
1 Introduction
In recent years, there has been much discussion around the Metaverse, an envisioned fully immersive iteration of the Internet, composed of a shared collection of interoperable 3D virtual worlds. The Metaverse remains a vision that is in its early stages Cheng et al. (2022), but many social virtual worlds usually associated with it thrived during the Covid-19 pandemic Kerdvibulvech (2022). In particular, interest surged in the wake of Facebook’s rebranding into Meta in late 2021, which signaled the company’s quest to realize such an ambitious vision Kraus et al. (2022). In these virtual worlds, interaction with the environment and with other people is done primarily through an avatar, an embodied representation of the user. Further, avatar customization (e.g., via virtual wearables) is a significant in-world factor of purchase intent, as well as one of the main pleasures for users Bleize & Antheunis (2019). However, it has been argued Zhou et al. (2018) that many of the ownership policies for in-world digital goods —such as those used in avatar customization— are detrimental to users due to the imbalance in terms of rights between platforms and users. For example, traditionally digital assets are not transferable independently of the platform; hence, these assets remain locked-in inside the virtual world or are forever lost when the platform ceases to function.
In response, blockchain technology has been heralded by many tech enthusiasts as a decentralized and transparent approach to digital property in virtual worlds Cannavo & Lamberti (2020). Arguably, Decentraland is the earliest and most popular blockchain-based virtual world, touted to be the first one “built, governed, and owned by its users” in the project’s official website. In Decentraland, users can transfer in-world land, usernames, and avatar wearables, either as part of a sale or for free. All of these transfers are recorded on the blockchain Ethereum, or in the case of wearables on Polygon, a more scalable and compatible blockchain. Hence, there is a trove of public data regarding the dynamics of this social virtual world.
Based on these data, Trujillo and Bacciu 2023 were among the first to conduct a quantitative study of Decentraland’s wearables. In their analysis, it emerged that a single account —related to the project Decentral Games (DG)— was by far the one with most published wearables. DG manages a few in-world virtual casinos, and possession of DG wearables is a requirement to earn money (in the form of cryptocurrency) by playing ICE Poker, the main project’s attraction (depicted in Figure 1). In addition, owners can delegate their DG wearables to other users, with whom earnings are split based on the wearable’s level, which can be also upgraded for a fee. Hence, DG wearables transcend the mere aesthetic in-world function of virtual garments, with ownership being enticing to users of Decentraland due to their earning potential.
Figure 1: Screenshot of an ICE Poker game session inside a Decentral Games casino on Decentraland. Note the avatar wearing the sleeveless shirt with the name of the game, which allows the player to earn cryptocurrency, as well as the croupier being based on the Internet meme Pepe the Frog.
Therefore, herein we delve into how DG and ICE Poker influence the dynamics of ownership transfers of wearables and in-world visits in Decentraland, a virtual world that touts its social and creative aspects of all kinds, but that seems to be driven by gambling. In particular, we seek to answer the following research questions (RQs):
RQ1:
Are transfers of Decentraland wearables disproportionately related to those created by Decentral Games?
RQ2:
Do Decentraland users visit significantly more ICE Poker casinos compared to other in-world locations?
To answer RQ1, we measure the share of wearables designed by DG in the overall dynamics of transferring ownership of Decetranland’s wearables in its different forms, including creation, sale, and destruction, as well as how the upgrading mechanism implemented by DG affects such transfers. To answer RQ2, we quantify the visits —both in time and number of visitors— made on parcels in which ICE Poker casinos are located, compared to the rest of the in-world land. We also consider and describe the time series characteristics of these visits on a daily basis.
1.1 Contributions and significance
Our main contribution is the first quantitative analysis regarding the influence that online gambling has had on Decentraland’s social virtual world. Thanks to its trailblazing development based on blockchain and tokenomics, Decentraland gives a glimpse of how new kinds of immersive social media might work on the so-called Metaverse. Therefore, studying the use (and potential abuse) of the platform’s affordances and economic incentives in the form of online gambling is important to better shape the future mechanisms and legal framework of social virtual worlds.
2 Background
In this section we briefly describe the main mechanisms, jargon, and events of both Decentraland and DG referred to in the related work, analyses, and discussion thereof.
2.1 Decentraland
Publicly released in February 2020, Decentraland’s social virtual world revolves around the so-called tokenomics, i.e., a decentralized economy based on the exchange of assets via cryptographic tokens on the blockchain Lo & Medda (2020). On Ethereum and Polygon, these exchanges are made from one address to another address via smart contracts, which set the automatic rules of the transactions. These addresses identify either an account or a smart contract that can participate on the blockchain.
Decentraland has its own cryptocurrency, called MANA, with which users can buy in-world digital assets, such as parcels of land and wearables. The proof of ownership of these assets is in the form of non-fungible tokens (NFTs) —unique and non-interchangeable, unlike the fungible cryptocurrencies. Possession of MANA is also used as the basis of voting power for platform decisions via the decentralized autonomous organization (DAO) of Decentraland, a form of blockchain-based governance without a central authority Goldberg & Schär (2023).
The world (see Figure 2) is composed of 90,601 square land parcels distributed on a 301$\times$301 grid, whose coordinates range in both axes from -150 to 150. Land is divided into two main types: public and privately owned. Public land is in turn divided into roads (10.4% of the world map) and plazas (4%), and cannot be sold. Before the public release of the project, 37.4% of the world map was allocated for districts, i.e., groups of adjacent parcels of privately owned land that share similar user interests. District parcels were assigned via an auctioning process to different groups, which respectively manage the land within their boundaries. Among the biggest districts we find, for example, the gambling-focused Vegas City and the shopping-oriented Fashion Street.
Figure 2: Decentraland world map, where the biggest districts are indicated in violet and roads in gray.Vegas City district is home to Diamond Hands City and The Stronghold (both highlighted in yellow), ICE Poker casinos owned by Decentral Games. They cover only a minuscule surface of the map, accounting for less than 0.1%.
Decentraland wearables are virtual garments, accessories or full body costumes (called skins) that can be used to clothe one’s avatar or completely change its appearance. Wearables are grouped into collections of one or more items. Each item represents a distinct design from which one or more NFTs can be minted, i.e, a new unique token ID representing the asset is created and assigned. Mints are registered on the blockchain as transfers from a special null address. In addition, a given item has a rarity limit (from a predefined list) that sets its maximum supply of tokens —e.g., a unique item can be minted only once, while a common item 100,000 times. To publish a wearable collection, creators pay a fee and submit their designs to the approval committee appointed by the DAO. Once approved, creators can sell their wearables via Decentraland’s Marketplace (for a fee that goes to the DAO), or they can mint and transfer the respective NFTs to a blockchain address (their own or not), either as a gift or to be sold on a different marketplace.
Originally, in the first version of wearables, these were created via a contest with approved designers and minted on Ethereum. However, this proved to be too limiting and prohibitive due to the then high transaction fees on that blockchain. Therefore, the wearables system was changed to the second and current version described above, now open to all and based on the less costly and more scalable Polygon blockchain. For the launch of both versions, a fashion event was held to showcase this creative endeavour and promote the newly published wearables; these events would be the precursors of more elaborate fashion initiatives. Arguably, the most important subsequent initiative was the first edition of the Metaverse Fashion Week in March 2022, a series of in-world social events that involved renowned fashion brands such as Tommy Hilfiger and Dolce & Gabbana.
2.2 Decentral Games
The project DG started in 2019, a year prior to Decentraland’s public release, and has played a part in many initiatives in the virtual world ever since. In fact, DG established one of the first in-world casinos, with games such as roulette or poker, and it was one of the few creators involved in the short-lived first version of Decentraland wearables. DG also established its own DAO, as well as its utility token and namesake DG, which grants access to rewards and voting power in the project’s governance.
In October 2021, DG released ICE Poker, touted to be a Metaverse play-to-earn game, i.e., one in which users are able to earn cryptocurrency simply by playing. In preparation, DG also launched the cryptocurrency ICE to function as the in-game currency for rewards and minting wearables. The following month, DG also published several exclusive ICE-themed wearables that granted attractive bonuses while playing the game, based on their ranking level. These were made available for minting in three batches to users that had a certain amount of DG and ICE.
Inside Decentraland, there are two DG virtual casinos dedicated to ICE Poker, both located in the Vegas District. The first one and more popular is called The Stronghold; the second one, Diamond Hands City, is exclusive to accounts that possess ICE-themed wearables with the highest rank. Outside Decentraland, DG wearables also give access to a related but separate mobile and web game released on September 2022, now called ICE Poker Arcade. In this version of the game, instead of cryptocurrency players earn electronic gadgets, DG merchandise or virtual assets such as wearables.
The level of access and/or reward provided by DG wearables varies by item, as illustrated by the examples in Figure 3; the more expensive the better. Once inside either game mode, players can wear up to five wearables to augment their rewards. For certain ICE-themed wearables, it is also possible to upgrade their ranking level (ranging from 1 to 5) to gain better rewards by paying a smaller fee compared to buying one. Additionally, all DG wearables can be delegated to other users, who can then play and split their earnings with the owner of the wearable, with the latter retaining the rights over the asset on the blockchain and being able of revoking the delegation at any moment.
Figure 3: Examples of wearables created by Decentral Games: a) Venetian Mask (ICE Rank 1), which gives 1–7% bonus in ICE Poker; b) DG Suited Tracksuit Bottom, which can be used to play ICE Poker Arcade Mode; and c) Paludamentum (ICE Rank 5), which gives 35–45% bonus in ICE Poker.
3 Related Work
While we are still in the early stages of Metaverse development, numerous independent virtual worlds have already sprung up over the past few decades. Notable examples include Second Life, Meta Horizon Worlds, and massively multiplayer online (MMO) games such as World of Warcraft, Minecraft, Roblox and Fortnite. However, these virtual environments have faced significant criticism regarding the governance and ownership of in-world assets. In Second Life, for instance, there has been a perceived emergence of a new form of virtual feudalism Grimmelmann (2009). World of Warcraft has grappled with conflicts between the game’s publisher and players over agency and ownership of in-game assets Glas (2010). Additionally, there has been a general separation between content and platform ownership in virtual worlds, often to the detriment of users, whose rights over bought or created in-platform content are usually restricted Zhou et al. (2018).
In response to these challenges, a wave of blockchain-based virtual experiences has emerged Cannavo & Lamberti (2020); Huynh-The et al. (2023). Virtual worlds such as Decentraland Ordano et al. (2017), The Sandbox Madrid et al. (2020), Axie Infinity Nguyen et al. (2021), Upland Honigstein et al. (2019), and Illuvium Warwick & Warwick (2023) all attempt to mitigate the aforementioned issues by leveraging the main strengths of blockchain technology (transparency, security, and decentralization) for various purposes. These include verifying the digital ownership of virtual assets, enabling cryptocurrency-based economies, and implementing decentralized governance systems.
In particular, Decentraland constitutes a standout example, being the earliest and most popular virtual world based on blockchain. It has reached a wide audience and gone through many changes over several years, hence it offers a trove of publicly available data for the study of emerging phenomena in these novel asset-oriented social media. However, existing research on Decentraland has primarily focused on landholding assets Dowling (2022); Goldberg et al. (2021); Guidi & Michienzi (2022) or examined the platform’s tokens within the broader context of the NFT ecosystem Nadini et al. (2021). There is a lack of studies regarding the other main kinds of digital assets available in Decentraland, namely usernames and wearables for avatars, with the latter having attracted the attention of the fashion industry Gonzalez (2022).
Fashion in social virtual worlds has been a subject of research for years, especially in Second Life, which has been active since 2003 and introduced the concept of customizable virtual spaces and avatars Bardzell et al. (2010). Despite criticisms that it did not live up to public expectations Rodriguez Sanchez & Garcia-Badell (2022), Second Life maintains a robust economy centered on avatar customization, with many users becoming in-world fashion designers, creating clothing and accessories for avatars Boellstorff (2015).
Recently, the fashion industry’s attention has shifted towards popular MMO games, particularly through the creation of branded skins—cosmetic options for characters, with an increasing interest in newer blockchain-based virtual worlds Rodriguez Sanchez & Garcia-Badell (2022). Notably, several prestigious luxury brands are venturing into NFT collectibles Joy et al. (2022), participating in events like the Metaverse Fashion Week or designing wearables for Decentraland Gonzalez (2022). Among the scarce scholarly investigation into Decentraland’s wearables, we find a quantitative analysis in which it emerged that they exhibit peculiar dynamics Trujillo & Bacciu (2023). For instance, wearables are mainly distributed for free to boost engagement in other cryptoasset or Metaverse initiatives. Only a minor portion of them are actually sold on the platform’s marketplace. Furthermore, the price of these wearables is largely influenced by their pre-established rarity, which escalates the cost even further for unique items (i.e., having a set minting limit of one). The study also identified a specific account that stands out in terms of wearable items collections, having published more than 80 at the time of the study. This is noteworthy, given that over half of all wearable creators in Decentraland typically author only a single collection. This creator is associated with Decentral Games, a company whose primary offering is a poker gaming experience within Decentraland virtual casinos.
Online gambling is a topic of research in its own right. This includes studies on the legality of online gambling and cryptocurrencies in different jurisdictions Zborowska et al. (2012); Yeoh (2017), the role of virtual communities in either motivating gamblers or protecting them from excessive behaviour Sirola et al. (2021); Savolainen et al. (2022), the overlap of personality and demographic traits among cryptocurrency traders and problem gamblers Johnson et al. (2023); Delfabbro et al. (2021), how those traits are partly different in professional gamblers Newall & Talberg (2023), and how playing habits in low-stakes gambling might influence the perception of financial risk in general Armstrong et al. (2018).
4 Methods
For our study, we focus on the two years after the launch of Decentraland wearables on Polygon, that is, from June 2021 to May 2023. We excluded the short-lived first version of wearables because these were already deprecated by the launch of ICE Poker. In the following, we describe the overall approach to answer our RQs and then we detail how we performed the respective data collection.
4.1 Overall approach
For RQ1 (wearable transfers) we first retrieved all wearables created during the period of study, marking those created by DG based on metadata from Decentraland. We then retrieved from the Polygon blockchain all the transfers related to all wearables, as well as their sales made in the most prominent marketplaces. We then analyzed the proportion between DG wearables and the rest in terms of the number of transfers as well as sales volume (both in number and USD). We also analyzed the share of transfers related to upgrades, which we retrieved from the respective log events on Polygon.
For RQ2 (in-world visits to DG poker casinos), we first collected in-world tracking data of users, available only from 12 July 2022. Hence, we analyzed the last 10 months of the two-year period of interest, in terms of the number of accounts and time spent visiting the parcels in which DG poker casinos are located, compared to the rest of Decentraland’s map.
Our descriptive and statistical analyses mostly concern proportions, i.e., the share of wearables or land associated with DG compared with the rest of Decentraland. For RQ1, this is either for the whole two-year period or on a monthly basis therein. For RQ2, this is either for a ten-month period or on a daily basis, as we have a narrower time frame but a finer time resolution for data of in-world visits. To test for significant differences in proportions we used two-sided tests with significance level $\alpha=0.05$. When analyzing samples we used $\chi^{2}$ tests, either goodness-of-fit test or association test according to the case; when analyzing populations we used a binomial test; for both kinds of test we use relative risk (RR) to measure the effect size. For daily in-world visits, we also used iterative STL (i.e., seasonal and trend decomposition using LOESS) to analyze the time series, including trend and seasonal strengths (from 0 to 1).
4.2 Data collection
For wearable metadata (e.g., collection, category, creator), we used several of the openly available Decentraland’s query endpoints on TheGraph, a service for querying blockchain networks, and whose definitions are open source. [^1] https://github.com/decentraland For transfers (including mints), we collected the respective transactions from Polygon via the blockchain web service Alchemy. For wearable sales, we collected data from two different sources: 1) Decentraland’s Marketplace for internal sales, using the aforementioned TheGraph query endpoints; and 2) OpenSea, the largest overall NFT marketplace and main means for external sales, as well as the initial default marketplace for DG wearables.
We also retrieved from Yahoo! Finance the price of the few different cryptocurrencies used for the collected wearable sales during the two-year period of study. To normalize the price of sales, we first converted it to USD, then we adjusted it for inflation on a monthly basis as of December 2023, based on the Consumer Price Index (CPI) by the Bureau of Labor Statistics of the United States. Price changes of these cryptocurrencies also serve to gauge the interest on the studied platforms. Indeed, as can be seen in Figure 4, the ICE token was released around Meta’s rebranding in November 2021, after the latter provoked a sudden price increase in Decentraland’s MANA (fourfold in just a couple of days). This was a considerable surge, even taking into account the already upward trend of MATIC in the wider Polygon ecosystem. However, the collapse of several interconnected cryptocurrencies the following May marked the onset of the 2022 cryptowinter OECD (2022), which affected a large share of blockchain projects and lasted until mid 2023.
Figure 4: Cryptocurrency price in USD during the two-year period of study for: ICE by Decentral Games, MANA by Decentraland, and MATIC by Polygon.
For data on DG wearable upgrades, we again used Alchemy to retrieve the log events for the period of study of DG’s smart contract which manages the respective on-chain transactions, called “iceRegistrant”. We then used the aforementioned wearable metadata to derive the ranking level and rarity limit of the upgraded tokens.
For in-world user tracking, we utilized the corresponding data archives from Atlas Data Corp, a blockchain analytics company that obtained a grant from Decentraland’s DAO to finance a system to keep record of user sessions. These data are log events (created every circa 20 seconds) of the current location of every connected user on the several servers that manage the 3D rendering of in-world scenes and avatars on Decentraland. Given that this system started operating in mid July 2022, we narrowed our analysis of in-world visits to only 10 months, from August 2022 to May 2023. Moreover, we take into consideration only users with an Ethereum-compatible address, that is, we discard the small minority of guest users not associated with an account. We then transformed the more than 677 million log events for this time frame into sessions with a start and end for each land parcel and account address. Finally, we segmented the respective sessions’ duration on a daily basis, adding 10 seconds to each to compensate for possible shortenings due to the 20-second polling rate of the log events.
5 Analyses and Results
For the two-year period of study, a total of 4,629 collections with 8,450 wearable items (i.e., designs) were published by 1,517 distinct creators. Of these, we identified and manually verified two accounts that created all of the DG wearables, based on marketplace offerings, item names, and user profiles on Decentraland. Both DG accounts (0.1% of creators) published 117 collections with 723 items (8.6% of the total available items). Henceforth, we refer to these as DG wearables.
5.1 RQ1: Overall wearable transfers
[…]
In our analysis on a monthly basis, we noticed great variation in the share of DG wearables over time, but interestingly, the variations are different for transfers and sales. On one hand, and as illustrated in Figure 5, the total volume of transfers reached its highest values in March and April of 2022, which coincides with the first edition of the Metaverse Fashion Week at the end of March. In those two months the share of DG wearables was only 3.3% and 4.5%, respectively. However, the transfers monthly share of DG wearables increased noticeably from September 2022 onward, reaching its peak in May 2023 with 85.3%, while the rest of wearable transfers decreased in number. This DG share increase is mostly related to a few wearable items with a common rarity limit and which give access only to ICE Poker Arcade, whose app was released precisely at the end of September. On the other hand, for the monthly sales volume in USD (illustrated in Figure 6), the highest monthly value for both total volume and share of DG wearables (62.9%) is November 2021, the month in which the first DG collections focused on ICE Poker were released for minting. Moreover, the timing also coincides with the wake of Meta’s rebranding, which provoked a surge and sustained high price in both ICE and MANA for that month, as shown in Figure 4.
Figure 5: Monthly number of transfers of Decentraland wearables on the Polygon blockchain and respective share associated with Decentral Games (DG). The upward trend in the bottom panel is related to wearables that grant access to the ICE Poker Arcade app, released in September 2022.
Figure 6: Monthly sales volume in USD of Decentraland wearables —made via the marketplaces of Decentraland and OpenSea— and the respective share associated with Decentral Games (DG). The bottom panel’s spike in November 2021 is related to the release of ICE-themed DG wearables and the high valuation of ICE in the wake of Meta’s rebranding.
For the two-year period there were 58,791 mints of wearables with an ICE ranking level, of which 42,390 (72.1%) were via upgrades. Each upgrade implies three wearable NFT transfers on Polygon: 1) transfer of the original NFT from the owner to the upgrading smart contract; 2) burning of this NFT by transferring it to a special dead address, effectively removing it from circulation; and 3) minting a new NFT for the owner from a higher-ranking wearable item within the same collection. Hence, there were 127,170 transfers related to upgrades, representing 14.2% of DG wearable transfers and 2.1% of all transfers. Interestingly, only 10% of the mints of ICE-ranking wearables with a level higher than 1 was not related to an upgrade. In other words, 90% of owners preferred to burn a lower-ranking token and mint a higher-ranking token via an upgrade, rather than buy (or wait to win) one without sacrificing the former.
5.2 RQ2: In-world visits of poker casinos
[…]
When analyzed on a daily basis, visits to the ICE Poker casinos are subject to relatively small fluctuations —compared to the rest of the world map— both in terms of time spent (see Figurge 7) and number of distinct visitor accounts (see Figure 8). For visits to ICE Poker casinos, we have a moderate but steady downward trend for both time spent and visitors, with a trend strength (based on STL decomposition) of 0.98 and 0.99, respectively. In the rest of the world map, the trend for time spent is slightly downward and with a strength of 0.83 for time spent, while the trend is visibly downward with a strength 0.91 for visitors. In other words, over time there were less distinct accounts visiting Decentraland, but these collectively spent more time on it. However, upon further inspection this increase in time spent per visitor is mostly due to a few accounts with a much higher number of hours than average. Indeed, in terms of share, time spent on ICE Poker casinos reached its highest levels at the beginning of September 2022, up to 53%, but it decreased to circa 20% in May 2023 due the the aforementioned outlier users in the rest of the world map. Nevertheless, for visitors, the share at the beginning of the ten-month period was of circa 32%, and at the end it increased to circa 41%. Curiously, the number of visitors of ICE Poker casinos had a modest weekly seasonality strength of 0.66, with the number of visitors slightly decreasing on weekends; for the other time series the seasonality strength was less than 0.5.
Figure 7: Stacked area chart of daily time spent (in hours) visiting Decentraland. ICE Poker casinos had a remarkable share, with a daily median of 20%. The highest overall spike was reached during the Metaverse Music Festival (MVMF) in mid-November.
Figure 8: Stacked area chart of daily distinct accounts that visited Decentraland. ICE Poker casinos had a striking share, with a daily median of 33%. The Metaverse Music Festival (MVMF) in mid-November had the highest spike overall, while the Metaverse Fashion Week (MVFW) at the end of March had the only spike during the last four months.
These trend and seasonality strength differences and shares by land category are in part explained by their dissimilar spikes and surges. On one hand, visits to the ICE Poker casinos had no spikes nor dips. There was, however, a short-lived surge (for both time and accounts) that started at the end of August 2022, when DG gifted special wearables to circa 3,000 of the most active players as a reward and in preparation for its new Tournament game mode. On the other hand, in the rest of the world map there were several visit spikes, surges, and dips. In particular, the Decentraland Metaverse Music Festival, held on 10–13 November 2022, was the event that provoked the most visits (both in time and accounts). The festival was free to attend, and its line up involved famous artists such as Björk, Ozzy Osbourne, Dillon Francis, and Soulja Boy. Incidentally, the Metaverse Fashion Week of 2023 also provoked a spike in visitors at the end of March, but not in time spent and only during a period of four months in which the number of distinct visitors had already remarkably declined.
6 Discussion
Based on our results, we can respond affirmatively to both RQ1 and RQ2: wearable transfers on Decentraland are indeed disproportionately related to those created by DG, and the ICE Poker casinos do receive significantly (and remarkably) more visits than other in-world locations. In the following, we first discuss our main findings regarding the influence of DG and ICE Poker on Decentraland, we comment on the in-world economic mechanisms that allowed such influence to exist, we reflect upon the ambiguous legal standing of ICE Poker, and, finally, we describe the main limitations of our study.
6.1 Main findings
Our findings confirm that DG and ICE Poker have great influence on Decentraland wearable and visit dynamics. Still, this influence has varied over time, on account of the initiatives that both Decentraland and DG have carried out to attract and retain users, and due to external market forces.
The Metaverse Fashion Week is arguably the most important initiative around Decentraland wearables. Both editions of 2022 and 2023 attracted well-established fashion brands, were featured in mainstream media, and piqued the interest of the general public. The initiative, however, did not meet expectations. Based on a qualitative study on the 2023 edition Miikkulainen-Gilbert (2023), most user complaints were related to the on-boarding complexity of setting up an Ethereum account, a lackluster in-world user experience, the context switch between in-world and traditional web-based interaction for many events, the technical limitations of wearable 3D modeling, and the need of paid approval even for wearables designed for oneself. In our analysis of visits we indeed detected a spike of distinct visitors during the 2023 edition at the end of March (see Figure 8), but there was no visible effect on the time spent on the platform (see Figure 7), and the number of visitors continued its downward trend afterward. The fact that in 2024 there was no Metaverse Fashion Week confirms the failure of the initiative to keep up the momentum going. On the other hand, the Metaverse Music Festival attracted much more visitors that spent considerable time in Decentraland (see Figures 7 and 8), but nevertheless visits did not visibly increase in the following weeks. Unlike in-world fashion events, music concerts have been present for many years in social virtual worlds and MMO games, and during the Covid-19 pandemic they were seen as an attractive alternative for the struggling live music industry Çeli̇k (2023). It remains to be seen if this Decentraland initiative will continue to gather interest in future editions.
As for DG, the main initiative after the release of ICE Poker wearables was the introduction of Tournament mode, which changed the rules to win prizes and claim ICE tokens. These changes were deemed necessary due to the quick pace with which delegates, who mostly did not own other kinds of DG assets (e.g., wearables or its utility token), withdrew ICE to convert it to other cryptocurrency or fiat money, which accelerated its inflation. Based on a financial report report by an external firm made in collaboration with DG, the aforementioned changes achieved their goal by reducing the inflation of ICE by circa 58% Sun (2023). In order to not alienate core delegates with such changes, DG gifted wearables to the 3,000 most active ones, which seems to explain the surge in visitors and time spent in ICE Poker casinos at the end of August 2022 (see Figures 7 and 8). Another important DG initiative was the launch of its Arcade mode and related web and mobile apps in September 2022, which allow users to play and earn non-monetary prizes without the need of visiting Decentraland; possession of DG wearables, however, is still necessary. This explains the increase in the share of transfers of DG wearables at the end of the period of study (see Figure 5) without the corresponding increase in visits share to ICE Poker casinos.
Concerning external market forces, the ICE token and the initial ICE Poker wearables were by happenstance released a few days shy of Meta’s rebranding. This image change of such a behemoth corporation is considered to be the main factor behind the surge in popularity of the term Metaverse and the hype thereof Dolata & Schwabe (2023). Consequently, the valuation of DG’s newly introduced cryptocurrency and wearables reached prominent heights, along with Decentraland’s own cryptocurrency, as can be seen in Figures 4 and 6.
Nevertheless, the interest on the Metaverse has now greatly subsided.
Over time, these platforms have tried to adapt to the hectic blockchain community, volatile cryptocurrencies, and stricter regulations.
Indeed, at the moment of writing (early 2024), DG is transitioning into a single utility and cryptocurrency token called BAG, to simplify its tokenomics and counter some of the remaining inflationary issues of ICE Decentral Games (2024). DG also plans to revamp its games, sunsetting the current system of upgrades and delegations in favor of improving the utility of owning a DG wearable, and it is also consolidating its operating license, which is currently on questionable legal standing.
6.2 On the legality of ICE Poker
After a few years of uncertainty regarding its legal status, online gambling is now banned or heavily regulated in most jurisdictions Zborowska et al. (2012). Similarly, for many years and due to their novelty, blockchain cryptocurrencies remained in a regulatory limbo in most jurisdictions Yeoh (2017). However, after many instances of scams and fraud —some very large and highly mediatized, as in the case of the exchange FTX— in the last few years there has been a push to outright ban or closely regulate cryptocurrencies Trautman et al. (2022). ICE Poker is a combination of both gambling and cryptocurrencies, hence, this begs the simple question on whether it is a legal operation or not. The answer is, however, murky.
Firstly, on the project’s front page it is stated that DG is operated by the company Kidderminster Cheltenham, whose official address is in Costa Rica. This is most likely due to the country’s laissez-faire approach for both online gambling and cryptocurrencies  Kerr (2018), its tax haven tendencies Baron (2019), and the fact that its operating licenses are often used to offer an online gambling service in foreign jurisdictions in which the licensing process is more rigid or non-existing Fischer (2021). Moreover, in the official project documentation it is stated that the company behind DG is called Web4 LTD, which is established in the British Virgin Islands, the place of jurisdiction regarding the terms and conditions of the project, and a British overseas territory known for being a tax haven and having links to dubious online gambling operations Macegoniuk (2021).
Curiously, in the project’s website there is also a legal opinion regarding ICE Poker by a consultant from the Singaporean firm Jacque Law Zhirong (2022), in which it is stated that the game “would not constitute ‘gambling’, ‘betting’, ‘lottery’ under the RGA [Singaporean Remote Gambling Act of 2014]” because “gambling laws are not intended to cover games whereby there is no staking or wagering of money or money’s worth”. Indeed, possession of a DG wearable is necessary to earn ICE, but the wearable has independent value outside of ICE Poker and its use is not exclusive to DG. In addition, users play with chips, which according to the opinion Zhirong (2022), “are non-transferable tokens that hold no value or functionality other than to keep score when players play ICE Poker”. Based on the number of chips and wearables they possess, players periodically receive ICE, which can be converted to fiat money, but on an external exchange such as Binance. This convoluted mechanism is similar to the one used offline by pachinko parlours in Japan to circumvent gambling laws Brooks et al. (2008): pachinko players win small steel balls, which can be exchanged for a receipt with a nominal value, which in turn is used to claim a low-valued physical item (e.g., a specific pack of washing powder), and finally, the player can exchange this item for a lesser nominal value in cash outside the parlour on a nearby “independent office”.
Of course, the above legal opinion applies only to Singapore. In the case of other jurisdictions DG is more cautious, claiming that the onus is on players. In fact, to participate in its virtual casinos, players must guarantee they are not residents of one of the non-supported regions listed on the project’s website. At the moment of writing (early 2024), there were 36 non-supported regions, including high-income countries such as the United States, the United Kingdom, France, Germany, and Australia. ICE Poker Arcade might nonetheless be supported in these regions, given that users play to win physical or digital prizes, not money or cryptocurrency. To address the issue in a more solid manner, recently DG acquired a 90% stake in Virtue Gaming Ltd, the entity that holds a Malta license to enable real-money gambling in Virtue Poker, the first licensed cryptocurrency-based online poker. DG intends to expand this coveted license to its virtual casinos and other games, which would remove doubts regarding their current legal status in many jurisdictions.
6.3 Limitations
In spite of our careful approach, several limiting factors could influence the interpretation and applicability of our results. In particular, despite having public access to all of the wearable transactions on Polygon, we had to make use of sample data for information not readily available in order to better comprehend both transfers and visits.
One such limitation concerns sales data, as we only collected data from two marketplaces. Tracking every single transfer that is also a sale is burdensome because each marketplace has its own smart contracts and sales protocol. Moreover, these might even change over time. For instance, we noticed a sales drop that lasted several days on OpenSea in May 2022, which upon further inspection resulted from a migration to their current protocol called SeaPort. Nonetheless, the marketplaces we selected should cover the vast majority of wearable sales, as Decentraland’s Marketplace is the official and default venue, and OpenSea is the largest NFT marketplace and was the initial default venue to sell wearables by DG, which now has its own marketplace.
Another potential limitation regards in-world tracking of users. We did not use the term active users in our analysis because we only considered the presence of a given account address in the log events, i.e., we did not filter out potential bots or users that could be deemed inactive. In line with other kinds of social media, bots are also a plague in social virtual worlds, with automated avatars having a noticeable impact on these Varvello & Voelker (2010). In addition, a user might be authentic (i.e, not a bot) but be away from keyboard while their avatar is still present in the virtual world. Determining an avatar who represents an “authentic active user”, however, depends on an arbitrary set of characteristics and mechanisms. Indeed, calculating the number of daily active users on Decentraland has sparked much public debate, as different definitions and mechanisms might result in vastly different numbers Thompson et al. (2022). Our approach, on the other hand, is relatively simplistic, but at the same time it better responds RQ2, as in any case these avatars —active or inactive, human or bot— occupy space and resources on Decentraland’s virtual world.
7 Conclusions
The concept of the Metaverse has in recent years sparked the collective imagination on novel and creative ways in which we could interact within immersive social virtual worlds and represent ourselves via our avatars. Decentraland —whose economy, digital assets, and governance are all based on blockchain technology— offers a glimpse on how a potential Metaverse based on tokenomics might work in the future. Indeed, Decentraland has been a pioneer in developing a decentralized social virtual world in which asset ownership is radically different from previous virtual worlds. Moreover, some of its economic and artistic initiatives with worldwide renowned personalities and companies, such as the Metaverse Fashion Week and the Metaverse Music Festival, have reached the mainstream media and gathered much public attention. At present, however, based on our results we can affirm that ICE Poker is the main driver in the dynamics of wearables and visits in Decentraland, having thus great influence on the platform in general.

### v2

@@
 Abstract
-Decentraland is a blockchain-based social virtual world touted to be a creative space owned by its community, unlike previous virtual worlds.
-Its users can create and publish wearables, virtual garments to customize avatars, which can be then sold or given away via the blockchain.
-Decentral Games (DG), a single project owning two prominent in-world casinos, has by far created the most wearables, with these being necessary to earn cryptocurrency in their flagship game ICE Poker.
-We thus present a comprehensive study that investigates how DG and ICE Poker influence the overall dynamics of Decentraland wearables and in-world visits.
+Decentraland is a blockchain-based social virtual world touted to be a creative space owned by its community.
+In it, users can publish wearables used to customize avatars, which can be then sold or given away via blockchain transfers.
+Decentral Games (DG), a single project owning prominent in-world casinos, has by far created the most wearables, necessary to earn cryptocurrency in its flagship game ICE Poker.
+Herein, we present a comprehensive study on how DG and ICE Poker influence the dynamics of wearables and in-world visits in Decentraland.
 To this end, we analyzed 5.9 million wearable transfers made on the Polygon blockchain (and related sales) over a two-year period, and 677 million log events of in-world user positions in an overlapping 10-month period.
-We found that the influence of DG and Ice Poker is not only significant, but also substantial for transfers and sales monetary value of wearables, and very large for daily unique visitors and time spent in the virtual world.
-Despite several alternative in-world economic and artistic initiatives in Decentraland, some of which have attracted much attention from the general public, online poker appears to be the main driver of the analyzed dynamics.
-Our work thus contributes to the current understanding of user behavior in social virtual worlds and it is among the first to study the emerging phenomenon of blockchain-based online gambling in virtual spaces.
+We found that the platform-wise number of transfers and sales monetary value of wearables were disproportionally related to DG, and that its two ICE Poker casinos (less than 0.1% of the world map) represented a very large average share of daily unique visitors (33%) and time spent in the virtual world (20%).
+Despite several alternative in-world economic and artistic initiatives in Decentraland, some of which have attracted much attention from the general public, a single third-party online poker game appears to be the main driver of the analyzed dynamics.
+Our work thus contributes to the current understanding of user behavior in social virtual worlds, and it is among the first to study the emerging phenomenon of blockchain-based online gambling in virtual reality spaces.
 keywords:
@@
 However, it has been argued Zhou et al.
-(2018) that many of the ownership policies for in-world digital goods —such as those used in avatar customization— are detrimental to users due to the imbalance in terms of rights between platforms and users.
-For example, traditionally digital assets are not transferable independently of the platform; hence, these assets remain locked-in inside the virtual world or are forever lost when the platform ceases to function.
+(2018) that many of the ownership policies for in-world digital goods —such as those used in avatar customization— are disadvantageous to users.
+This is primarily due to the imbalance in rights between the users and the proprietors of the virtual world platforms.
+For example, traditionally digital assets are not transferable independently of the platform; hence, these assets remain locked-in inside the virtual world or are forever lost when the underlying platform ceases to function.
 In response, blockchain technology has been heralded by many tech enthusiasts as a decentralized and transparent approach to digital property in virtual worlds Cannavo & Lamberti (2020).
@@
 In Decentraland, users can transfer in-world land, usernames, and avatar wearables, either as part of a sale or for free.
-All of these transfers are recorded on the blockchain Ethereum, or in the case of wearables on Polygon, a more scalable and compatible blockchain.
+All of these transfers are recorded on the blockchain Ethereum, or in the case of wearables on Polygon, a compatible and more scalable blockchain.
 Hence, there is a trove of public data regarding the dynamics of this social virtual world.
@@
 Note the avatar wearing the sleeveless shirt with the name of the game, which allows the player to earn cryptocurrency, as well as the croupier being based on the Internet meme Pepe the Frog.
-Therefore, herein we delve into how DG and ICE Poker influence the dynamics of ownership transfers of wearables and in-world visits in Decentraland, a virtual world that touts its social and creative aspects of all kinds, but that seems to be driven by gambling.
+Therefore, intrigued by both the peculiar behavior of the aforementioned account managed by DG and the economic incentive to play in its casinos, herein we delve into how DG and ICE Poker influence the dynamics of ownership transfers of wearables and in-world visits in Decentraland.
 In particular, we seek to answer the following research questions (RQs):
@@
 2 Background
-In this section we briefly describe the main mechanisms, jargon, and events of both Decentraland and DG referred to in the related work, analyses, and discussion thereof.
+In this section we briefly describe the main mechanisms, jargon, and events of both Decentraland and Decentral Games referred to in the related work, analyses, and discussion thereof.
+2.1 Blockchain technologies
+Blockchains are decentralized, immutable ledgers that facilitate secure and transparent recording of transactions across a network of computers.
+They are the technological backbone of cryptocurrencies such as Bitcoin (BTCNakamoto (2008) and Ether (ETHWoodet al.
+(2014), digital currencies that enable financial transactions without the need for a central authority like a bank.
+Blockchains also provide the underlying mechanisms for non-fungible tokens (NFTs), unique identifiers often used to represent ownership of digital objects (e.g, art and collectibles) Nadini et al.
+(2021).
+Unlike fungible cryptocurrencies, NFTs are unique and non-interchangeable.
+Furthermore, blockchains enable decentralized autonomous organizations (DAOs), a form of token-based governance without a central authority, allowing for distributed decision-making and management Hassan & De Filippi (2021).
+The Ethereum blockchain Woodet al.
+(2014), in particular, is a versatile platform that supports all of the aforementioned applications, utilizing smart contracts —self-executing programs that automatically enforce the terms of an agreement.
+Decentraland’s features are based on the Ethereum blockchain, except for transactions concerning wearables; these are instead based on Polygon Kanani et al.
+(2021), a higher layer scaling solution with its own cryptocurrency (MATIC); it is compatible with Ethereum and is designed to improve transaction speeds and reduce costs while maintaining security and decentralization.
 2.2 Decentraland
 Publicly released in February 2020, Decentraland’s social virtual world revolves around the so-called tokenomics, i.e., a decentralized economy based on the exchange of assets via cryptographic tokens on the blockchain Lo & Medda (2020).
-On Ethereum and Polygon, these exchanges are made from one address to another address via smart contracts, which set the automatic rules of the transactions.
+On Ethereum and Polygon, these exchanges are made from one address to another address via the aforementioned smart contracts, which set the automatic rules of the transactions.
 These addresses identify either an account or a smart contract that can participate on the blockchain.
 Decentraland has its own cryptocurrency, called MANA, with which users can buy in-world digital assets, such as parcels of land and wearables.
-The proof of ownership of these assets is in the form of non-fungible tokens (NFTs) —unique and non-interchangeable, unlike the fungible cryptocurrencies.
-Possession of MANA is also used as the basis of voting power for platform decisions via the decentralized autonomous organization (DAO) of Decentraland, a form of blockchain-based governance without a central authority Goldberg & Schär (2023).
+The proof of ownership of these assets is in the form of NFTs.
+Possession of MANA is also used as the basis of voting power for platform decisions via the decentralized autonomous organization of Decentraland Goldberg & Schär (2023).
 The world (see Figure 2) is composed of 90,601 square land parcels distributed on a 301$\times$301 grid, whose coordinates range in both axes from -150 to 150.
@@
 DG also established its own DAO, as well as its utility token and namesake DG, which grants access to rewards and voting power in the project’s governance.
-In October 2021, DG released ICE Poker, touted to be a Metaverse play-to-earn game, i.e., one in which users are able to earn cryptocurrency simply by playing.
+In October 2021, DG released ICE Poker, touted to be a Metaverse Play-to-Earn (P2E) game, i.e., one in which users are able to earn cryptocurrency simply by playing.
 In preparation, DG also launched the cryptocurrency ICE to function as the in-game currency for rewards and minting wearables.
@@
 However, these virtual environments have faced significant criticism regarding the governance and ownership of in-world assets.
-In Second Life, for instance, there has been a perceived emergence of a new form of virtual feudalism Grimmelmann (2009).
+In Second Life, for instance, there has been a perceived emergence of a new form of virtual feudalism Grimmelmann (2009), while Roblox has been criticised for profiting from child labor Parkin (2022) and for allowing harmful content and games with gambling-like mechanisms in its user-generated virtual worlds Kou & Gui (2023).
 World of Warcraft has grappled with conflicts between the game’s publisher and players over agency and ownership of in-game assets Glas (2010).
@@
 There is a lack of studies regarding the other main kinds of digital assets available in Decentraland, namely usernames and wearables for avatars, with the latter having attracted the attention of the fashion industry Gonzalez (2022).
+3.1 Fashion in Virtual Worlds
 Fashion in social virtual worlds has been a subject of research for years, especially in Second Life, which has been active since 2003 and introduced the concept of customizable virtual spaces and avatars Bardzell et al.
@@
 Notably, several prestigious luxury brands are venturing into NFT collectibles Joy et al.
-(2022), participating in events like the Metaverse Fashion Week or designing wearables for Decentraland Gonzalez (2022).
-Among the scarce scholarly investigation into Decentraland’s wearables, we find a quantitative analysis in which it emerged that they exhibit peculiar dynamics Trujillo & Bacciu (2023).
-For instance, wearables are mainly distributed for free to boost engagement in other cryptoasset or Metaverse initiatives.
-Only a minor portion of them are actually sold on the platform’s marketplace.
-Furthermore, the price of these wearables is largely influenced by their pre-established rarity, which escalates the cost even further for unique items (i.e., having a set minting limit of one).
-The study also identified a specific account that stands out in terms of wearable items collections, having published more than 80 at the time of the study.
-This is noteworthy, given that over half of all wearable creators in Decentraland typically author only a single collection.
-This creator is associated with Decentral Games, a company whose primary offering is a poker gaming experience within Decentraland virtual casinos.
-Online gambling is a topic of research in its own right.
-This includes studies on the legality of online gambling and cryptocurrencies in different jurisdictions Zborowska et al.
-(2012); Yeoh (2017), the role of virtual communities in either motivating gamblers or protecting them from excessive behaviour Sirola et al.
+(2022), participating in events such as the Metaverse Fashion Week or designing wearables for Decentraland Gonzalez (2022), as well as retailing and engaging in other kinds of customer engagement within Decentraland’s social virtual world Goldberg et al.
+(2023).
+In fact, Decentraland’s wearables are mainly distributed for free to boost engagement in other cryptoasset or Metaverse initiatives, with only a minor portion of them (3.4%) being sold on the platform’s marketplace Trujillo & Bacciu (2023).
+Besides wearables, other kinds of NFTs available within the virtual world (e.g., tickets, memberships) are likewise transferred for free as part of promotional activities instead of being part of a sale transaction Goldberg et al.
+(2023).
+Perhaps unsurprisingly, when sold in Decentraland’s marketplace, the price of wearables is largely influenced by their pre-established rarity Trujillo & Bacciu (2023), which escalates even further for unique items —i.e., having a set minting limit of one.
+Incidentally, over half of all wearable creators in Decentraland typically author only a single collection (mostly with a single item), but as mentioned in the Introduction, a specific account related to Decentral Games stands out having published circa 80 collections with more than 600 items total in a year and a half of activity Trujillo & Bacciu (2023).
+These wearables are the only way to get access to the games played on the company’s casinos within Decentraland, with the primary offering being ICE Poker, a virtual poker gaming experience.
+3.2 Poker and Online Gambling
+The chance-based nature of the game of poker appeals to a wide audience of casual players, while the element of skill, particularly evident over many hands, entices professional players Fiedler & Rock (2009).
+The introduction of Internet to gambling activities has provided easy access for the general public, round-the-clock access, shorter intervals between bets, and instant reinforcements Griffiths & Barnes (2008).
+After 2003, online poker experienced tremendous growth, with a total market value reaching 3 to 4 billion USD by 2010, with approximately 5.5 million players participating in virtual poker rooms across the globe Fiedler & Wilcke (2011).
+Among these players, a subset has emerged as professionals, especially adept at multitabling —simultaneously playing multiple games to further reduce the element of chance.
+As reliance on poker earnings grew, these players developed a strong extrinsic motivation to play, even if not entirely removed from the gameplay.
+Further, it has been found that playing for real money strongly influences both game dynamics and player experiences in poker games Zaman et al.
+(2014).
+Notably, professional gamblers share psychological traits with problem gamblers, highlighting the complex interplay between skill and risk Newall & Talberg (2023).
+Problem gamblers, often drawn to low-stakes gambling, may develop distorted risk perceptions and ultimately underestimate financial risks in general Armstrong et al.
+(2018).
+Interestingly, players are socially motivated to gamble by the virtual communities they form, while these same communities also serve as safeguards against excessive behavior Sirola et al.
 (2021); Savolainen et al.
-(2022), the overlap of personality and demographic traits among cryptocurrency traders and problem gamblers Johnson et al.
+(2022).
+Online poker’s unique subculture —which includes websites, forums, and chat rooms from all around the world— fosters cooperative behavior among players through the sharing of information and strategies.
+Nonetheless, the game’s inherent competitiveness persists and is reflected in the hierarchical structures within these global communities O’Leary & Carroll (2013).
+Despite online gambling being a global phenomenon, its legality varies across jurisdictions, thereby posing significant challenges for both players and operators as they attempt to determine whether a specific game is legal to play or not Zborowska et al.
+(2012).
+The advent of blockchain technology has ignited enthusiasm due to its potential to improve fairness and transparency in the sector, promising decentralized and autonomous solutions for regulators and operators alike.
+However, at the same time it ultimately adds complexity to the legal framework of online gambling, due in great part to the current regulatory uncertainties of cryptocurrencies and other blockchain-based assets Gainsbury & Blaszczynski (2017).
+3.3 Crypto Asset Trading
+Cryptocurrency trading has emerged as a disruptive force in the financial landscape since the inception of Bitcoin in 2008.
+While market capitalizations have reached trillions of dollars, the trading environment remains characterized by high volatility, often leading to substantial crashes followed by significant rebounds Fang et al.
+(2022); OECD (2022).
+NFTs entered the scene in 2014, gaining widespread attention in early 2021 due to very high-profile sales.
+Digital art, music albums, and even tweets and sports highlights have been registered as unique crypto assets and subsequently sold for millions of dollars each, causing considerable hype Houser & Holden (2022).
+NFT users exhibit a profound enthusiasm for the foundational technologies that facilitate unique avenues for creative expression and the development of innovative content creation business models.
+However, these positive sentiments are counterbalanced by the contentious matter of digital content ownership, the prevalence of low-quality NFTs, the risk of scams, potential money laundering activities, and the evolving framework of legal measures Sharma et al.
+(2022).
+The literature also shows an overlap of personality and demographic traits among problem gamblers and cryptocurrency traders Johnson et al.
 (2023); Delfabbro et al.
-(2021), how those traits are partly different in professional gamblers Newall & Talberg (2023), and how playing habits in low-stakes gambling might influence the perception of financial risk in general Armstrong et al.
+(2021b).
+In both cryptocurrency trading and gambling, decisions are often based on limited information, short-term motives for gain, and with high uncertainty; hence, people attracted to latter are likely to be attracted to the former Delfabbro et al.
+(2021a).
+Numerous operators in the crypto gambling sector provide access to a variety of betting options, games, and online casinos.
+However, these platforms often pose a significant risk as they are easily accessible to minors and vulnerable individuals, and typically offer minimal to no consumer protection features.
+The legal landscape for these platforms is also a concern, as they remain largely unregulated in many countries Andrade et al.
 (2023).
+Gambling aside, the legality of trading both cryptocurrencies and NFTs also varies greatly across jurisdictions.
+Some countries embrace these digital assets, while others impose strict regulations.
+Factors include tax treatment, anti-money laundering compliance, and investor protection.
+Researchers and policymakers continue to grapple with these complexities, seeking a balance between industry innovation and safeguarding the public Yeoh (2017).
+3.4 Play-to-Earn Games
+The emergence of blockchain technologies has also given rise to a new form of monetized online gaming known as “Play-to-Earn” (P2E) games.
+In P2E games, players receive monetary rewards for their gameplay, mainly in the form of the game’s specific cryptocurrency.
+Axie Infinity —a Pokémon-style battle game with collectible axolotl-like creatures— is arguably one of the most well-known examples, making use of cryptocurrencies and NFTs for its in-game economy.
+Its popularity peaked in 2021, particularly in the Philippines and later Venezuela, where it represented a viable source of earnings for low-income people during the height of the pandemic De Jesus et al.
+(2022).
+However, extrinsically motivated gameplay, i.e., mainly driven by external rewards such as financial gain, can reduce intrinsic motivation and consequently have negative implications for player experience Delic & Delfabbro (2024); Hong et al.
+(2023).
+A survey conducted with a group of Axie Infinity players in 2020–21 has shown that, while some players enjoyed social interactions, overall intrinsic motivation tended to be low, causing high amounts of stress and unwanted tension to reach daily quotas De Jesus et al.
+(2022).
+This assessment was later corroborated by a study that analyzed online conversations about the game Delic & Delfabbro (2024).
+Interestingly, part of the initial success of Axie Infinity was due to delegation models that emerged organically within the community, called scholarships De Jesus et al.
+(2022).
+These models allow players to share their knowledge or delegate their gameplay (and assets) to others, in exchange for a share of the rewards.
+Nonetheless, scholarship models and related delegation mechanisms within P2E games are potentially exploitative.
+While P2E games offer players from less-advantaged countries the prospect of earning income above the national average wage, players might find themselves participating in a scheme that provides low-cost labor for wealthier actors who hold valuable NFTs Delic & Delfabbro (2024).
+Curiously, ICE Poker integrates by design some elements of P2E games, offering the possibility to earn tokens just by participating in certain events, and a delegation mechanism similar to the community-based scholarships in Axie Infinity.
 4 Methods
@@
 Figure 4: Cryptocurrency price in USD during the two-year period of study for: ICE by Decentral Games, MANA by Decentraland, and MATIC by Polygon.
-For data on DG wearable upgrades, we again used Alchemy to retrieve the log events for the period of study of DG’s smart contract which manages the respective on-chain transactions, called “iceRegistrant”.
+For data on DG wearable upgrades, we again used Alchemy to retrieve the log events for the period of study of DG’s smart contract on Polygon [^2] Address 0xC9a67eD1472A76d064C826B54c144Ca00DAE4015 which manages the respective on-chain transactions, called “iceRegistrant”.
 We then used the aforementioned wearable metadata to derive the ranking level and rarity limit of the upgraded tokens.
-For in-world user tracking, we utilized the corresponding data archives from Atlas Data Corp, a blockchain analytics company that obtained a grant from Decentraland’s DAO to finance a system to keep record of user sessions.
+For in-world user tracking, we utilized the corresponding data archives from Atlas Data Corp, a blockchain analytics company that obtained a grant  [^3] https://decentraland.org/governance/project/?id=7422a99e-8a25-4625-9b30-ab688de5dade from Decentraland’s DAO to finance a system to keep record of user sessions.
 These data are log events (created every circa 20 seconds) of the current location of every connected user on the several servers that manage the 3D rendering of in-world scenes and avatars on Decentraland.
@@
 5 Analyses and Results
-For the two-year period of study, a total of 4,629 collections with 8,450 wearable items (i.e., designs) were published by 1,517 distinct creators.
-Of these, we identified and manually verified two accounts that created all of the DG wearables, based on marketplace offerings, item names, and user profiles on Decentraland.
+5.1 RQ1: Overall wearable transfers
+From June 2021 to May 2023, a total of 4,629 collections with 8,450 wearable items (i.e., designs) were published by 1,517 distinct creators.
+Of these, we identified and manually verified two accounts [^4] Addresses 0x17a253c2ac0d5ba92cadbbf665e3390c9913dc5d and 0x00e5d44f6a296c10f159486f842838bd68f13e32 that created all of the DG wearables, based on marketplace offerings, item names, and user profiles on Decentraland.
 Both DG accounts (0.1% of creators) published 117 collections with 723 items (8.6% of the total available items).
 Henceforth, we refer to these as DG wearables.
-5.1 RQ1: Overall wearable transfers
 […]
@@
 Based on our results, we can respond affirmatively to both RQ1 and RQ2: wearable transfers on Decentraland are indeed disproportionately related to those created by DG, and the ICE Poker casinos do receive significantly (and remarkably) more visits than other in-world locations.
-In the following, we first discuss our main findings regarding the influence of DG and ICE Poker on Decentraland, we comment on the in-world economic mechanisms that allowed such influence to exist, we reflect upon the ambiguous legal standing of ICE Poker, and, finally, we describe the main limitations of our study.
+In the following, we first discuss our main findings concerning the influence of DG and ICE Poker on Decentraland, we comment on the initiatives and events affecting such influence over time, we reflect upon the ambiguous legal standing of ICE Poker, we explore the social implications of our findings, and, finally, we describe the main limitations of our study.
 6.1 Main findings
 Our findings confirm that DG and ICE Poker have great influence on Decentraland wearable and visit dynamics.
-Still, this influence has varied over time, on account of the initiatives that both Decentraland and DG have carried out to attract and retain users, and due to external market forces.
+Just two DG accounts out of 1,517 wearable creators published 8.6% of the total available items, with these being twice more likely to be minted (i.e., involved in a creation transfer) based on the maximum supply limit compared to the rest, particularly at the end of the period of study.
+DG wearables also had a monetary sales volume an order of magnitude higher despite the considerably fewer sales.
+Moreover, in spite of covering less that 0.1% of the land, ICE Poker casinos had a striking share of time spent and unique visitors in Decentraland, with a respective daily median of 20% and 33%.
+Further, the vast majority of these visits were done to a single casino: The Stronghold.
+Still, these dynamics have varied over time, on account of the initiatives that both Decentraland and DG have carried out to attract and retain users, and due to external market forces.
 The Metaverse Fashion Week is arguably the most important initiative around Decentraland wearables.
@@
 In order to not alienate core delegates with such changes, DG gifted wearables to the 3,000 most active ones, which seems to explain the surge in visitors and time spent in ICE Poker casinos at the end of August 2022 (see Figures 7 and 8).
-Another important DG initiative was the launch of its Arcade mode and related web and mobile apps in September 2022, which allow users to play and earn non-monetary prizes without the need of visiting Decentraland; possession of DG wearables, however, is still necessary.
+Another important DG initiative was the launch of its Arcade mode and related web and mobile apps in September 2022, which allow users to play and earn non-monetary prizes without the need of visiting Decentraland; possession of DG wearables is still necessary, however.
 This explains the increase in the share of transfers of DG wearables at the end of the period of study (see Figure 5) without the corresponding increase in visits share to ICE Poker casinos.
@@
 Consequently, the valuation of DG’s newly introduced cryptocurrency and wearables reached prominent heights, along with Decentraland’s own cryptocurrency, as can be seen in Figures 4 and 6.
-Nevertheless, the interest on the Metaverse has now greatly subsided.
-Over time, these platforms have tried to adapt to the hectic blockchain community, volatile cryptocurrencies, and stricter regulations.
+Nevertheless, the initial hype regarding NFTs has subsided (and to a lesser extent also that of the Metaverse), with attention having shifted to generative artificial intelligence, which came to the spotlight with the release of ChatGPT Echauri (2024).
+In addition, platforms based on tokenomics must continually adapt to the hectic blockchain community, volatile cryptocurrencies, and stricter regulations OECD (2022).
 Indeed, at the moment of writing (early 2024), DG is transitioning into a single utility and cryptocurrency token called BAG, to simplify its tokenomics and counter some of the remaining inflationary issues of ICE Decentral Games (2024).
@@
 DG intends to expand this coveted license to its virtual casinos and other games, which would remove doubts regarding their current legal status in many jurisdictions.
+6.3 Social implications
+Unlike most previous social virtual worlds and social media in general, platform transparency and openness are stated to be paramount in Decentraland.
+The project strives toward these qualities by means of its use of public records on the blockchain, user-oriented ownership of in-world assets, public and open-source code repositories, governance via a DAO, and initiatives to overtly monitor the ecosystem —e.g., by financing the collection of in-world tracking data we used to analyze visits.
+These characteristics, common to many blockchain-based projects, also offer unprecedented opportunities for both industry and academia to develop and understand novel sociotechnical systems, in the form of virtual reality spaces but not only Cannavo & Lamberti (2020).
+Nevertheless, the current levels of transparency and openness in Decentraland do not preclude the abuse of the platform’s mechanisms.
+For instance, in June 2022 DG submitted a proposal to Decentraland’s DAO for a grant, [^5] https://decentraland.org/governance/proposal/?id=99c66b90-e2d2-11ec-9000-175d8dd584b8 asking the equivalent to one million USD in MANA to increase its liquidity pool for ICE rewards.
+The grant was approved by 56% of the voting power, which is based on ownership of Decentraland MANA, usernames, and land holdings.
+DG —one of the earliest and most prominent actors in the platform— accounted for almost a third of the voting power in its own favor.
+This is an illustrative example of how a decentralized governance is not necessarily a neutral one.
+Rather, in its current form Decentraland’s DAO gives way to automated situations reminiscent of the principal–agent problem Goldberg & Schär (2023), because “what is best for a dominant individual is not necessarily a good choice for the virtual world as a whole as well as the majority of its stakeholders”.
+Other examples of potential platform abuse are the aforementioned drastic reward changes introduced with Tournament mode and the current revamping of ICE Poker itself.
+These highlight that despite the assurances of public transactions and smart contracts, most of the functionality of the poker system —and the virtual world itself— lives outside the blockchain, and is thus subject to arbitrary changes by its controlling entity.
+In addition, a third-party actor was able to exploit the affordances provided by the virtual world, which lead to deviations in user behavior unforeseen by the platform’s creators.
+The decision by DG to use Decentraland wearables as access tokens to play in its online casino games has significantly influenced the overall acquisition trends of wearables at the platform level, tilting the motivating factors for acquiring wearables from intrinsic (e.g, the aesthetic enjoyment of the garment) to extrinsic (e.g., the earning utility of the token).
+Hence, functional wearables might be appreciated solely for their utility, whereas cosmetic ones might be deemed less valuable due to their absence of utility, which weakens Decentraland’s claim to be a creative social space.
+Besides, based on our results in-world visits seem be greatly influenced by the earning potential within ICE Poker casinos, instead of the ludic enjoyment that these and other world areas offer.
+However, an overabundance of extrinsic motivation can often undermine an experience by eroding the spontaneous intrinsic motivation, especially for games Delic & Delfabbro (2024); Hong et al.
+(2023).
+At the same time, in online poker the earning prospects are partially integrated into the gameplay, thus becoming part of the intrinsic motivation to play, but with the possibility of leading to problem gambling without the proper containment mechanisms Zaman et al.
+(2014).
+ICE Poker sits at the intersection of gamblers and cryptoasset traders, with potentially troubling consequences as both profiles might exhibit similar self-destructive behavior, such as engaging in risk-taking and thrill-seeking activities, and a susceptibility to addiction Johnson et al.
+(2023); Delfabbro et al.
+(2021b).
+We should not forget, however, that in Ice Poker, as with other play-to-earn games, a considerable amount of grunt play is done by low-income people who do not own assets and for whom game enjoyment is secondary.
+It is a new form of digital exploitation —where the lion’s share goes to the owner of the NFT who lends it to the less fortunate— and which could lead to increased social disparity and potentially exacerbate economic inequalities Zaucha (2024).
+For all the above reasons, the study of the emerging phenomena in novel social platforms such as Decentraland and games such as Ice Poker is increasingly important.
+A better understanding of their underlying mechanisms and dynamics will allow, for instance, polycymakers to be better informed and prepared to enact the necessary changes to tackle these and other related issues that might emerge in the next few years.
 6.4 Limitations
@@
 Decentraland —whose economy, digital assets, and governance are all based on blockchain technology— offers a glimpse on how a potential Metaverse based on tokenomics might work in the future.
-Indeed, Decentraland has been a pioneer in developing a decentralized social virtual world in which asset ownership is radically different from previous virtual worlds.
+The platform has been a pioneer in developing a decentralized social virtual world in which asset ownership is radically different from previous virtual worlds.
 Moreover, some of its economic and artistic initiatives with worldwide renowned personalities and companies, such as the Metaverse Fashion Week and the Metaverse Music Festival, have reached the mainstream media and gathered much public attention.
-At present, however, based on our results we can affirm that ICE Poker is the main driver in the dynamics of wearables and visits in Decentraland, having thus great influence on the platform in general.
+At present, however, based on our results we can affirm that ICE Poker by Decentral Games (with its enticing money-earning promise) is the main driver in the dynamics of wearables and visits in Decentraland, having thus great influence on the platform in general.
+This predominance of a single third-party online gambling initiative —which has transformed itself several times due to technical, economic, and legal uncertainties— raises doubts on if and how Decentraland’s current bet on tokenomics to realize its vision of the Metaverse will ultimately pay off.

### v3

@@
 Abstract
-Decentraland is a blockchain-based social virtual world touted to be a creative space owned by its community.
-In it, users can publish wearables used to customize avatars, which can be then sold or given away via blockchain transfers.
-Decentral Games (DG), a single project owning prominent in-world casinos, has by far created the most wearables, necessary to earn cryptocurrency in its flagship game ICE Poker.
-Herein, we present a comprehensive study on how DG and ICE Poker influence the dynamics of wearables and in-world visits in Decentraland.
+Decentraland is a blockchain-based social virtual world where users can publish and sell wearables to customize avatars.
+In it, the third-party Decentral Games (DG) allows players of its flagship game ICE Poker to earn cryptocurrency only if they possess certain wearables.
+Herein, we present a comprehensive study on how DG and its game influence the dynamics of wearables and in-world visits in Decentraland.
 To this end, we analyzed 5.9 million wearable transfers made on the Polygon blockchain (and related sales) over a two-year period, and 677 million log events of in-world user positions in an overlapping 10-month period.
-We found that the platform-wise number of transfers and sales monetary value of wearables were disproportionally related to DG, and that its two ICE Poker casinos (less than 0.1% of the world map) represented a very large average share of daily unique visitors (33%) and time spent in the virtual world (20%).
-Despite several alternative in-world economic and artistic initiatives in Decentraland, some of which have attracted much attention from the general public, a single third-party online poker game appears to be the main driver of the analyzed dynamics.
-Our work thus contributes to the current understanding of user behavior in social virtual worlds, and it is among the first to study the emerging phenomenon of blockchain-based online gambling in virtual reality spaces.
-keywords:
-social virtual worlds, Metaverse, online gambling, NFT, blockchain
-1 Introduction
+We found that these activities are disproportionately related to DG, with its ICE Poker casinos (less than 0.1% of the world map) representing a remarkable average share of daily unique visitors (33%) and time spent in the virtual world (20%).
+Despite several alternative initiatives within Decentraland, ICE Poker appears to drive user activity on the platform.
+Our work thus contributes to the understanding of how play-to-earn games influence user behavior in social virtual worlds, and it is among the first to study the emerging phenomenon of virtual blockchain-based gambling.
+Amaury TRUJILLO
+Istituto di Informatica e Telematica, Consiglio Nazionale delle Ricerche (IIT-CNR)
+Pisa, Italy
+amaury.trujillo@iit.cnr.it
+[Uncaptioned image] $0$-0001-6227-0944
+(Corresponding Author)
+Clara BACCIU
+Istituto di Informatica e Telematica, Consiglio Nazionale delle Ricerche (IIT-CNR)
+Pisa, Italy
+clara.bacciu@iit.cnr.it
+[Uncaptioned image] $0$-0002-8711-7812
+Matteo ABRATE
+Istituto di Informatica e Telematica, Consiglio Nazionale delle Ricerche (IIT-CNR)
+Pisa, Italy
+matteo.abrate@iit.cnr.it
+[Uncaptioned image] 0009-0004-6809-3570
+Keywords— social virtual worlds, play-to-earn, online gambling
+I.   Introduction
 In recent years, there has been much discussion around the Metaverse, an envisioned fully immersive iteration of the Internet, composed of a shared collection of interoperable 3D virtual worlds.
-The Metaverse remains a vision that is in its early stages Cheng et al.
-(2022), but many social virtual worlds usually associated with it thrived during the Covid-19 pandemic Kerdvibulvech (2022).
-In particular, interest surged in the wake of Facebook’s rebranding into Meta in late 2021, which signaled the company’s quest to realize such an ambitious vision Kraus et al.
-(2022).
-In these virtual worlds, interaction with the environment and with other people is done primarily through an avatar, an embodied representation of the user.
-Further, avatar customization (e.g., via virtual wearables) is a significant in-world factor of purchase intent, as well as one of the main pleasures for users Bleize & Antheunis (2019).
-However, it has been argued Zhou et al.
-(2018) that many of the ownership policies for in-world digital goods —such as those used in avatar customization— are disadvantageous to users.
-This is primarily due to the imbalance in rights between the users and the proprietors of the virtual world platforms.
-For example, traditionally digital assets are not transferable independently of the platform; hence, these assets remain locked-in inside the virtual world or are forever lost when the underlying platform ceases to function.
-In response, blockchain technology has been heralded by many tech enthusiasts as a decentralized and transparent approach to digital property in virtual worlds Cannavo & Lamberti (2020).
-Arguably, Decentraland is the earliest and most popular blockchain-based virtual world, touted to be the first one “built, governed, and owned by its users” in the project’s official website.
+The Metaverse remains a vision in its early stages, but many social virtual worlds usually associated with it thrived during the Covid-19 pandemic [33].
+In these worlds, interaction with the environment and with other people is done primarily through an avatar, an embodied representation of the user, with avatar customization being a significant in-world factor of purchase intent and one of the main pleasures for users [4].
+However, it has been argued [58] that many of the ownership policies for in-world digital goods —such as those used in avatar customization— are disadvantageous to users.
+This is primarily due to the imbalance in rights between the users and platform proprietors.
+Traditionally, digital assets are not transferable independently of the platform; hence, these remain locked inside the virtual world or are forever lost when the underlying platform ceases to function.
+In response, blockchain technology has been heralded by many tech enthusiasts as a decentralized and transparent approach to digital property in virtual worlds [9].
+Arguably, Decentraland is the earliest and most popular blockchain-based virtual world, touted to be the first one “built, governed, and owned by its users” in its official website.
 In Decentraland, users can transfer in-world land, usernames, and avatar wearables, either as part of a sale or for free.
-All of these transfers are recorded on the blockchain Ethereum, or in the case of wearables on Polygon, a compatible and more scalable blockchain.
-Hence, there is a trove of public data regarding the dynamics of this social virtual world.
-Based on these data, Trujillo and Bacciu 2023 were among the first to conduct a quantitative study of Decentraland’s wearables.
-In their analysis, it emerged that a single account —related to the project Decentral Games (DG)— was by far the one with most published wearables.
-DG manages a few in-world virtual casinos, and possession of DG wearables is a requirement to earn money (in the form of cryptocurrency) by playing ICE Poker, the main project’s attraction (depicted in Figure 1).
-In addition, owners can delegate their DG wearables to other users, with whom earnings are split based on the wearable’s level, which can be also upgraded for a fee.
-Hence, DG wearables transcend the mere aesthetic in-world function of virtual garments, with ownership being enticing to users of Decentraland due to their earning potential.
-Figure 1: Screenshot of an ICE Poker game session inside a Decentral Games casino on Decentraland.
-Note the avatar wearing the sleeveless shirt with the name of the game, which allows the player to earn cryptocurrency, as well as the croupier being based on the Internet meme Pepe the Frog.
-Therefore, intrigued by both the peculiar behavior of the aforementioned account managed by DG and the economic incentive to play in its casinos, herein we delve into how DG and ICE Poker influence the dynamics of ownership transfers of wearables and in-world visits in Decentraland.
+All these transfers are recorded on the Ethereum blockchain, or, in the case of wearables, on Polygon, a compatible and more scalable blockchain.
+Based on these data, Trujillo and Bacciu [51] were among the first to conduct a quantitative study of Decentraland’s wearables.
+They noticed that a single account —related to the project Decentral Games (DG)— had by far the most published wearables.
+DG manages a few in-world casinos, and possession of DG wearables is required to earn money (in the form of cryptocurrency) by playing ICE Poker, its main attraction (depicted in Figure I).
+Owners can also delegate their DG wearables to others, with whom earnings are split based on the wearable’s level.
+Hence, DG wearables transcend the mere aesthetic in-world function of virtual garments, with ownership being enticing due to their earning potential.
+Herein we thus delve into how DG and ICE Poker influence the dynamics of wearables’ ownership and in-world visits in Decentraland.
 In particular, we seek to answer the following research questions (RQs):
@@
 Do Decentraland users visit significantly more ICE Poker casinos compared to other in-world locations?
-To answer RQ1, we measure the share of wearables designed by DG in the overall dynamics of transferring ownership of Decetranland’s wearables in its different forms, including creation, sale, and destruction, as well as how the upgrading mechanism implemented by DG affects such transfers.
-To answer RQ2, we quantify the visits —both in time and number of visitors— made on parcels in which ICE Poker casinos are located, compared to the rest of the in-world land.
-We also consider and describe the time series characteristics of these visits on a daily basis.
-1.1 Contributions and significance
-Our main contribution is the first quantitative analysis regarding the influence that online gambling has had on Decentraland’s social virtual world.
-Thanks to its trailblazing development based on blockchain and tokenomics, Decentraland gives a glimpse of how new kinds of immersive social media might work on the so-called Metaverse.
-Therefore, studying the use (and potential abuse) of the platform’s affordances and economic incentives in the form of online gambling is important to better shape the future mechanisms and legal framework of social virtual worlds.
-2 Background
-In this section we briefly describe the main mechanisms, jargon, and events of both Decentraland and Decentral Games referred to in the related work, analyses, and discussion thereof.
-2.1 Blockchain technologies
+To answer RQ1, we measure the share of wearables designed by DG in the overall dynamics of ownership transfers in different forms, including creation, sale, and destruction, as well as how the upgrading mechanism implemented by DG affects such transfers.
+To answer RQ2, we analyze the visits —both in time spent and number of visitors— made on parcels in which ICE Poker casinos are located, compared to the rest of the in-world land.
+Figure I: Screenshot of an ICE Poker game session inside a Decentral Games casino.
+Our main contribution is therefore a quantitative analysis regarding the influence that online gambling has had on Decentraland’s social virtual world.
+In fact, integrating earning opportunities into virtual spaces has already resulted in serious repercussions in the real world[26, 41], emphasizing how virtual worlds can be both digital (in their essence) and real (in their effects) [7].
+Thus, studying the use (and potential abuse) of Decentraland’s affordances and economic incentives in the form of online gambling is important to better shape the future mechanisms and legal framework of social virtual worlds and the Metaverse in general.
+II.   Background
+In this section we briefly describe the main mechanisms, jargon, and events of both Decentraland and Decentral Games referred to in subsequent sections.
+A.   Blockchain technologies
 Blockchains are decentralized, immutable ledgers that facilitate secure and transparent recording of transactions across a network of computers.
-They are the technological backbone of cryptocurrencies such as Bitcoin (BTCNakamoto (2008) and Ether (ETHWoodet al.
-(2014), digital currencies that enable financial transactions without the need for a central authority like a bank.
-Blockchains also provide the underlying mechanisms for non-fungible tokens (NFTs), unique identifiers often used to represent ownership of digital objects (e.g, art and collectibles) Nadini et al.
-(2021).
+Decentraland is based on Ethereum [53], a versatile blockhain that enables smart contracts —self-executing programs that automatically enforce the terms of an agreement.
+These provide the underlying infrastructure for non-fungible tokens (NFTs), unique identifiers often used to represent ownership of digital objects (e.g, art and collectibles).
 Unlike fungible cryptocurrencies, NFTs are unique and non-interchangeable.
-Furthermore, blockchains enable decentralized autonomous organizations (DAOs), a form of token-based governance without a central authority, allowing for distributed decision-making and management Hassan & De Filippi (2021).
-The Ethereum blockchain Woodet al.
-(2014), in particular, is a versatile platform that supports all of the aforementioned applications, utilizing smart contracts —self-executing programs that automatically enforce the terms of an agreement.
-Decentraland’s features are based on the Ethereum blockchain, except for transactions concerning wearables; these are instead based on Polygon Kanani et al.
-(2021), a higher layer scaling solution with its own cryptocurrency (MATIC); it is compatible with Ethereum and is designed to improve transaction speeds and reduce costs while maintaining security and decentralization.
-2.2 Decentraland
-Publicly released in February 2020, Decentraland’s social virtual world revolves around the so-called tokenomics, i.e., a decentralized economy based on the exchange of assets via cryptographic tokens on the blockchain Lo & Medda (2020).
-On Ethereum and Polygon, these exchanges are made from one address to another address via the aforementioned smart contracts, which set the automatic rules of the transactions.
-These addresses identify either an account or a smart contract that can participate on the blockchain.
-Decentraland has its own cryptocurrency, called MANA, with which users can buy in-world digital assets, such as parcels of land and wearables.
-The proof of ownership of these assets is in the form of NFTs.
-Possession of MANA is also used as the basis of voting power for platform decisions via the decentralized autonomous organization of Decentraland Goldberg & Schär (2023).
-The world (see Figure 2) is composed of 90,601 square land parcels distributed on a 301$\times$301 grid, whose coordinates range in both axes from -150 to 150.
-Land is divided into two main types: public and privately owned.
-Public land is in turn divided into roads (10.4% of the world map) and plazas (4%), and cannot be sold.
+Smart contracts also enable decentralized autonomous organizations (DAOs), a form of token-based governance without a central authority, allowing for distributed decision-making and management [28].
+B.   Decentraland
+Publicly released in February 2020, Decentraland’s social virtual world revolves around the so-called tokenomics, i.e., an economy based on the exchange of assets via cryptographic tokens on blockchains [36].
+Decentraland has its own cryptocurrency, MANA, with which users can buy in-world NFTs for digital assets such as land parcels and wearables.
+Possession of MANA is also used as the basis of voting power for platform decisions via Decentraland’s DAO [22].
+The world (see Figure II) is composed of 90,601 square land parcels distributed on a 301$\times$301 grid.
 Before the public release of the project, 37.4% of the world map was allocated for districts, i.e., groups of adjacent parcels of privately owned land that share similar user interests.
-District parcels were assigned via an auctioning process to different groups, which respectively manage the land within their boundaries.
 Among the biggest districts we find, for example, the gambling-focused Vegas City and the shopping-oriented Fashion Street.
-Figure 2: Decentraland world map, where the biggest districts are indicated in violet and roads in gray.Vegas City district is home to Diamond Hands City and The Stronghold (both highlighted in yellow), ICE Poker casinos owned by Decentral Games.
-They cover only a minuscule surface of the map, accounting for less than 0.1%.
-Decentraland wearables are virtual garments, accessories or full body costumes (called skins) that can be used to clothe one’s avatar or completely change its appearance.
+Figure II: Decentraland world map, where the biggest districts are indicated in violet and roads in gray.
+District Vegas City is home to Diamond Hands City and The Stronghold (both highlighted in yellow), ICE Poker casinos owned by Decentral Games.
+They cover only a minuscule surface of the map: less than 0.1%.
+Decentraland wearables are virtual garments, accessories or full body costumes (called skins) used to change the appearance of one’s avatar.
 Wearables are grouped into collections of one or more items.
 Each item represents a distinct design from which one or more NFTs can be minted, i.e, a new unique token ID representing the asset is created and assigned.
-Mints are registered on the blockchain as transfers from a special null address.
 In addition, a given item has a rarity limit (from a predefined list) that sets its maximum supply of tokens —e.g., a unique item can be minted only once, while a common item 100,000 times.
@@
 Once approved, creators can sell their wearables via Decentraland’s Marketplace (for a fee that goes to the DAO), or they can mint and transfer the respective NFTs to a blockchain address (their own or not), either as a gift or to be sold on a different marketplace.
-Originally, in the first version of wearables, these were created via a contest with approved designers and minted on Ethereum.
-However, this proved to be too limiting and prohibitive due to the then high transaction fees on that blockchain.
-Therefore, the wearables system was changed to the second and current version described above, now open to all and based on the less costly and more scalable Polygon blockchain.
-For the launch of both versions, a fashion event was held to showcase this creative endeavour and promote the newly published wearables; these events would be the precursors of more elaborate fashion initiatives.
-Arguably, the most important subsequent initiative was the first edition of the Metaverse Fashion Week in March 2022, a series of in-world social events that involved renowned fashion brands such as Tommy Hilfiger and Dolce & Gabbana.
-2.3 Decentral Games
-The project DG started in 2019, a year prior to Decentraland’s public release, and has played a part in many initiatives in the virtual world ever since.
-In fact, DG established one of the first in-world casinos, with games such as roulette or poker, and it was one of the few creators involved in the short-lived first version of Decentraland wearables.
+C.   Decentral Games
+The project DG started in 2019 and has played a part in many initiatives in Decentraland ever since.
+In fact, DG established one of the first in-world casinos, with games such as roulette or poker, and it was a prominent creator of wearables since the beginning.
 DG also established its own DAO, as well as its utility token and namesake DG, which grants access to rewards and voting power in the project’s governance.
@@
 The following month, DG also published several exclusive ICE-themed wearables that granted attractive bonuses while playing the game, based on their ranking level.
-These were made available for minting in three batches to users that had a certain amount of DG and ICE.
-Inside Decentraland, there are two DG virtual casinos dedicated to ICE Poker, both located in the Vegas District.
-The first one and more popular is called The Stronghold; the second one, Diamond Hands City, is exclusive to accounts that possess ICE-themed wearables with the highest rank.
-Outside Decentraland, DG wearables also give access to a related but separate mobile and web game released on September 2022, now called ICE Poker Arcade.
-In this version of the game, instead of cryptocurrency players earn electronic gadgets, DG merchandise or virtual assets such as wearables.
-The level of access and/or reward provided by DG wearables varies by item, as illustrated by the examples in Figure 3; the more expensive the better.
-Once inside either game mode, players can wear up to five wearables to augment their rewards.
+Inside Decentraland, there are two DG virtual casinos dedicated to ICE Poker, both located in the Vegas City district.
+The first and more popular is called The Stronghold; the second, Diamond Hands City, is exclusive to ICE-themed wearables with the highest rank.
+Outside Decentraland, DG wearables also give access to a related but separate mobile and web game released on September 2022, called ICE Poker Arcade.
+In this version of the game, instead of cryptocurrency players earn electronic gadgets, DG merchandise or other virtual assets.
+The level of access and/or reward provided by DG wearables varies by item, as illustrated by the examples in Figure III; the more expensive the better.
+Within either game mode, players can wear up to five wearables to augment their rewards.
 For certain ICE-themed wearables, it is also possible to upgrade their ranking level (ranging from 1 to 5) to gain better rewards by paying a smaller fee compared to buying one.
-Additionally, all DG wearables can be delegated to other users, who can then play and split their earnings with the owner of the wearable, with the latter retaining the rights over the asset on the blockchain and being able of revoking the delegation at any moment.
-Figure 3: Examples of wearables created by Decentral Games: a) Venetian Mask (ICE Rank 1), which gives 1–7% bonus in ICE Poker; b) DG Suited Tracksuit Bottom, which can be used to play ICE Poker Arcade Mode; and c) Paludamentum (ICE Rank 5), which gives 35–45% bonus in ICE Poker.
-3 Related Work
-While we are still in the early stages of Metaverse development, numerous independent virtual worlds have already sprung up over the past few decades.
+Additionally, all DG wearables can be delegated to other players, who can then split their earnings with the wearable’s owner, with the latter retaining the rights over the asset on the blockchain and being able of revoking the delegation at any moment.
+Figure III: Examples of wearables created by Decentral Games: a) Venetian Mask (ICE Rank 1), which gives 1–7% bonus; b) DG Suited Tracksuit Bottom, which can be used to play Arcade Mode; and c) Paludamentum (ICE Rank 5), which gives 35–45% bonus.
+III.   Related Work
+While we are still in the early development stages of the Metaverse, numerous independent virtual worlds have already sprung up over the past few decades.
 Notable examples include Second Life, Meta Horizon Worlds, and massively multiplayer online (MMO) games such as World of Warcraft, Minecraft, Roblox and Fortnite.
 However, these virtual environments have faced significant criticism regarding the governance and ownership of in-world assets.
-In Second Life, for instance, there has been a perceived emergence of a new form of virtual feudalism Grimmelmann (2009), while Roblox has been criticised for profiting from child labor Parkin (2022) and for allowing harmful content and games with gambling-like mechanisms in its user-generated virtual worlds Kou & Gui (2023).
-World of Warcraft has grappled with conflicts between the game’s publisher and players over agency and ownership of in-game assets Glas (2010).
-Additionally, there has been a general separation between content and platform ownership in virtual worlds, often to the detriment of users, whose rights over bought or created in-platform content are usually restricted Zhou et al.
-(2018).
-In response to these challenges, a wave of blockchain-based virtual experiences has emerged Cannavo & Lamberti (2020); Huynh-The et al.
-(2023).
-Virtual worlds such as Decentraland Ordano et al.
-(2017), The Sandbox Madrid et al.
-(2020), Axie Infinity Nguyen et al.
-(2021), Upland Honigstein et al.
-(2019), and Illuvium Warwick & Warwick (2023) all attempt to mitigate the aforementioned issues by leveraging the main strengths of blockchain technology (transparency, security, and decentralization) for various purposes.
-These include verifying the digital ownership of virtual assets, enabling cryptocurrency-based economies, and implementing decentralized governance systems.
-In particular, Decentraland constitutes a standout example, being the earliest and most popular virtual world based on blockchain.
-It has reached a wide audience and gone through many changes over several years, hence it offers a trove of publicly available data for the study of emerging phenomena in these novel asset-oriented social media.
-However, existing research on Decentraland has primarily focused on landholding assets Dowling (2022); Goldberg et al.
-(2021); Guidi & Michienzi (2022) or examined the platform’s tokens within the broader context of the NFT ecosystem Nadini et al.
-(2021).
-There is a lack of studies regarding the other main kinds of digital assets available in Decentraland, namely usernames and wearables for avatars, with the latter having attracted the attention of the fashion industry Gonzalez (2022).
-3.1 Fashion in Virtual Worlds
-Fashion in social virtual worlds has been a subject of research for years, especially in Second Life, which has been active since 2003 and introduced the concept of customizable virtual spaces and avatars Bardzell et al.
-(2010).
-Despite criticisms that it did not live up to public expectations Rodriguez Sanchez & Garcia-Badell (2022), Second Life maintains a robust economy centered on avatar customization, with many users becoming in-world fashion designers, creating clothing and accessories for avatars Boellstorff (2015).
-Recently, the fashion industry’s attention has shifted towards popular MMO games, particularly through the creation of branded skins—cosmetic options for characters, with an increasing interest in newer blockchain-based virtual worlds Rodriguez Sanchez & Garcia-Badell (2022).
-Notably, several prestigious luxury brands are venturing into NFT collectibles Joy et al.
-(2022), participating in events such as the Metaverse Fashion Week or designing wearables for Decentraland Gonzalez (2022), as well as retailing and engaging in other kinds of customer engagement within Decentraland’s social virtual world Goldberg et al.
-(2023).
-In fact, Decentraland’s wearables are mainly distributed for free to boost engagement in other cryptoasset or Metaverse initiatives, with only a minor portion of them (3.4%) being sold on the platform’s marketplace Trujillo & Bacciu (2023).
-Besides wearables, other kinds of NFTs available within the virtual world (e.g., tickets, memberships) are likewise transferred for free as part of promotional activities instead of being part of a sale transaction Goldberg et al.
-(2023).
-Perhaps unsurprisingly, when sold in Decentraland’s marketplace, the price of wearables is largely influenced by their pre-established rarity Trujillo & Bacciu (2023), which escalates even further for unique items —i.e., having a set minting limit of one.
-Incidentally, over half of all wearable creators in Decentraland typically author only a single collection (mostly with a single item), but as mentioned in the Introduction, a specific account related to Decentral Games stands out having published circa 80 collections with more than 600 items total in a year and a half of activity Trujillo & Bacciu (2023).
-These wearables are the only way to get access to the games played on the company’s casinos within Decentraland, with the primary offering being ICE Poker, a virtual poker gaming experience.
-3.2 Poker and Online Gambling
-The chance-based nature of the game of poker appeals to a wide audience of casual players, while the element of skill, particularly evident over many hands, entices professional players Fiedler & Rock (2009).
-The introduction of Internet to gambling activities has provided easy access for the general public, round-the-clock access, shorter intervals between bets, and instant reinforcements Griffiths & Barnes (2008).
-After 2003, online poker experienced tremendous growth, with a total market value reaching 3 to 4 billion USD by 2010, with approximately 5.5 million players participating in virtual poker rooms across the globe Fiedler & Wilcke (2011).
-Among these players, a subset has emerged as professionals, especially adept at multitabling —simultaneously playing multiple games to further reduce the element of chance.
+In Second Life, for instance, there has been a perceived emergence of virtual feudalism [26].
+Roblox has been criticized for profiting from child labor [45] and for allowing harmful content and games with gambling-like mechanisms in its user-generated virtual worlds [35].
+World of Warcraft has also grappled with conflicts between the game’s publisher and players over agency and ownership of in-game assets [20].
+More generally, there has been a separation between content and platform ownership in virtual worlds, often to the detriment of users, whose rights over bought or created in-platform content are usually restricted [58].
+In response, a wave of blockchain-based platforms has emerged [9].
+Virtual worlds such as Decentraland, The Sandbox, Axie Infinity, and Upland, among others, attempt to mitigate the aforementioned issues by leveraging blockchain’s transparency, security, and decentralization for digital ownership, cryptocurrency transactions, and governance.
+Decentraland [43] is a standout example, being the earliest and most popular virtual world based on blockchain.
+It has reached a wide audience and gone through many changes over several years, hence it offers a trove of publicly available data for the study of emerging phenomena on these novel asset-oriented platforms.
+However, existing research on Decentraland has primarily focused on landholding assets [14, 21, 27] or examined the platform’s tokens within the broader context of the NFT ecosystem [39].
+There is a lack of studies regarding other kinds of digital assets available in Decentraland, namely avatar wearables, with the latter having attracted the attention of the fashion industry [24].
+A.   Fashion in Virtual Worlds
+Fashion in social virtual worlds has been a subject of research for years, especially in Second Life, which has been active since 2003 and introduced the concept of customizable virtual spaces and avatars [3].
+Despite its recent decline, Second Life maintained for years a robust economy centered on avatar customization, with many users becoming in-world fashion designers [6, 8].
+Recently, the fashion industry’s attention has shifted towards popular MMO games, particularly through the creation of branded skins, with an increasing interest in newer blockchain-based virtual worlds [46].
+Notably, several prestigious luxury brands are venturing into NFT collectibles [32], participating in events such as the Metaverse Fashion Week or designing wearables for Decentraland [24], as well as retailing and engaging in other kinds of customer engagement within its social virtual world [23].
+In fact, Decentraland’s wearables are mainly distributed for free to boost engagement in other cryptoasset or Metaverse initiatives such as virtual casinos, with only a minor portion of them (3.4%) being sold on the platform’s marketplace [51].
+B.   Poker and Online Gambling
+The chance-based nature of the game of poker appeals to a wide audience of casual players, while the element of skill, particularly evident over many hands, entices professional players [18].
+The introduction of the Internet to gambling activities has provided easy access for the general public, round-the-clock access, shorter intervals between bets, and instant reinforcements [25].
+After 2003, online poker experienced tremendous growth, with a total market value reaching 3 to 4 billion USD by 2010, with approximately 5.5 million players participating in virtual poker rooms across the globe [17].
+Among such players, a subset has emerged as professionals, especially adept at multitabling —simultaneously playing multiple games to further reduce the element of chance.
 As reliance on poker earnings grew, these players developed a strong extrinsic motivation to play, even if not entirely removed from the gameplay.
-Further, it has been found that playing for real money strongly influences both game dynamics and player experiences in poker games Zaman et al.
-(2014).
-Notably, professional gamblers share psychological traits with problem gamblers, highlighting the complex interplay between skill and risk Newall & Talberg (2023).
-Problem gamblers, often drawn to low-stakes gambling, may develop distorted risk perceptions and ultimately underestimate financial risks in general Armstrong et al.
-(2018).
-Interestingly, players are socially motivated to gamble by the virtual communities they form, while these same communities also serve as safeguards against excessive behavior Sirola et al.
-(2021); Savolainen et al.
-(2022).
+Further, it has been found that playing for real money strongly influences both game dynamics and player experiences in poker games [55].
+Notably, professional gamblers share psychological traits with problem gamblers, highlighting the complex interplay between skill and risk [40].
+Problem gamblers, often drawn to low-stakes gambling, may develop distorted risk perceptions and ultimately underestimate financial risks in general [2].
+Interestingly, players are socially motivated to gamble by the virtual communities they form, while these same communities also serve as safeguards against excessive behavior [49, 47].
 Online poker’s unique subculture —which includes websites, forums, and chat rooms from all around the world— fosters cooperative behavior among players through the sharing of information and strategies.
-Nonetheless, the game’s inherent competitiveness persists and is reflected in the hierarchical structures within these global communities O’Leary & Carroll (2013).
-Despite online gambling being a global phenomenon, its legality varies across jurisdictions, thereby posing significant challenges for both players and operators as they attempt to determine whether a specific game is legal to play or not Zborowska et al.
-(2012).
-The advent of blockchain technology has ignited enthusiasm due to its potential to improve fairness and transparency in the sector, promising decentralized and autonomous solutions for regulators and operators alike.
-However, at the same time it ultimately adds complexity to the legal framework of online gambling, due in great part to the current regulatory uncertainties of cryptocurrencies and other blockchain-based assets Gainsbury & Blaszczynski (2017).
-3.3 Crypto Asset Trading
-Cryptocurrency trading has emerged as a disruptive force in the financial landscape since the inception of Bitcoin in 2008.
-While market capitalizations have reached trillions of dollars, the trading environment remains characterized by high volatility, often leading to substantial crashes followed by significant rebounds Fang et al.
-(2022); OECD (2022).
-NFTs entered the scene in 2014, gaining widespread attention in early 2021 due to very high-profile sales.
-Digital art, music albums, and even tweets and sports highlights have been registered as unique crypto assets and subsequently sold for millions of dollars each, causing considerable hype Houser & Holden (2022).
-NFT users exhibit a profound enthusiasm for the foundational technologies that facilitate unique avenues for creative expression and the development of innovative content creation business models.
-However, these positive sentiments are counterbalanced by the contentious matter of digital content ownership, the prevalence of low-quality NFTs, the risk of scams, potential money laundering activities, and the evolving framework of legal measures Sharma et al.
-(2022).
-The literature also shows an overlap of personality and demographic traits among problem gamblers and cryptocurrency traders Johnson et al.
-(2023); Delfabbro et al.
-(2021b).
-In both cryptocurrency trading and gambling, decisions are often based on limited information, short-term motives for gain, and with high uncertainty; hence, people attracted to latter are likely to be attracted to the former Delfabbro et al.
-(2021a).
+Nonetheless, the game’s inherent competitiveness persists and is reflected in the hierarchical structures within these global communities [44].
+C.   Crypto Asset Trading
+Cryptocurrency has emerged as a disruptive force in the financial landscape since Bitcoin’s introduction in 2008.
+Nonetheless, its trading remains characterized by high volatility, often leading to substantial crashes followed by significant rebounds [16, 42].
+NFTs entered the scene in 2014, gaining widespread attention in early 2021 due to very high-profile sales and causing considerable hype [30].
+NFT buyers exhibit a profound enthusiasm for the foundational technologies that facilitate unique avenues for creative expression and the development of innovative content creation business models.
+However, these positive sentiments are counterbalanced by the contentiousness of digital ownership, prevalence of low-quality NFTs, risk of scams, potential money laundering activities, and evolving legal framework [48].
+There is also an overlap of personality and demographic traits among problem gamblers and cryptocurrency traders [31].
+In both cryptocurrency trading and gambling, decisions are often based on limited information, short-term motives for gain, and with high uncertainty; hence, people attracted to the latter are likely to be attracted to the former [11].
 Numerous operators in the crypto gambling sector provide access to a variety of betting options, games, and online casinos.
-However, these platforms often pose a significant risk as they are easily accessible to minors and vulnerable individuals, and typically offer minimal to no consumer protection features.
-The legal landscape for these platforms is also a concern, as they remain largely unregulated in many countries Andrade et al.
-(2023).
-Gambling aside, the legality of trading both cryptocurrencies and NFTs also varies greatly across jurisdictions.
-Some countries embrace these digital assets, while others impose strict regulations.
-Factors include tax treatment, anti-money laundering compliance, and investor protection.
-Researchers and policymakers continue to grapple with these complexities, seeking a balance between industry innovation and safeguarding the public Yeoh (2017).
-3.4 Play-to-Earn Games
-The emergence of blockchain technologies has also given rise to a new form of monetized online gaming known as “Play-to-Earn” (P2E) games.
-In P2E games, players receive monetary rewards for their gameplay, mainly in the form of the game’s specific cryptocurrency.
-Axie Infinity —a Pokémon-style battle game with collectible axolotl-like creatures— is arguably one of the most well-known examples, making use of cryptocurrencies and NFTs for its in-game economy.
-Its popularity peaked in 2021, particularly in the Philippines and later Venezuela, where it represented a viable source of earnings for low-income people during the height of the pandemic De Jesus et al.
-(2022).
-However, extrinsically motivated gameplay, i.e., mainly driven by external rewards such as financial gain, can reduce intrinsic motivation and consequently have negative implications for player experience Delic & Delfabbro (2024); Hong et al.
-(2023).
-A survey conducted with a group of Axie Infinity players in 2020–21 has shown that, while some players enjoyed social interactions, overall intrinsic motivation tended to be low, causing high amounts of stress and unwanted tension to reach daily quotas De Jesus et al.
-(2022).
-This assessment was later corroborated by a study that analyzed online conversations about the game Delic & Delfabbro (2024).
-Interestingly, part of the initial success of Axie Infinity was due to delegation models that emerged organically within the community, called scholarships De Jesus et al.
-(2022).
-These models allow players to share their knowledge or delegate their gameplay (and assets) to others, in exchange for a share of the rewards.
-Nonetheless, scholarship models and related delegation mechanisms within P2E games are potentially exploitative.
-While P2E games offer players from less-advantaged countries the prospect of earning income above the national average wage, players might find themselves participating in a scheme that provides low-cost labor for wealthier actors who hold valuable NFTs Delic & Delfabbro (2024).
-Curiously, ICE Poker integrates by design some elements of P2E games, offering the possibility to earn tokens just by participating in certain events, and a delegation mechanism similar to the community-based scholarships in Axie Infinity.
-4 Methods
+However, these platforms often pose a significant risk as they are easily accessible to minors and vulnerable individuals, and typically offer minimal to no consumer protection.
+Further, they still remain largely unregulated in many countries [1].
+D.   Play-to-Earn Games
+Blockchain technology has also boosted a form of online gaming known as “Play-to-Earn” (P2E), where players receive monetary rewards for their gameplay, mainly in the form of the game’s own cryptocurrency.
+Axie Infinity —a Pokémon-style battle game with collectible axolotl-like creatures and in-game tokenomics— is arguably one of the most well-known examples.
+Its popularity peaked in 2021, particularly in the Philippines and later Venezuela, where it represented a viable source of earnings for low-income people during the height of the Covid-19 pandemic [10].
+However, it is known that extrinsically motivated gameplay, i.e., mainly driven by external rewards such as financial gain, may reduce intrinsic motivation and consequently have negative implications for player experience [12, 29].
+In fact, a survey conducted with a group of Axie Infinity players in 2020–21 has shown that, while some players enjoyed social interactions, overall intrinsic motivation tended to be low, causing high amounts of stress and unwanted tension to reach daily quotas [10].
+This assessment was later corroborated by a study that analyzed online conversations about the game [12].
+Interestingly, part of the initial success of Axie Infinity was due to delegation models that emerged organically within the community, called scholarships [10].
+These allow players to share their knowledge or delegate their gameplay (and assets) to others for a share of the rewards.
+Nonetheless, while P2E games offer players from less-advantaged countries earning prospects above the national average wage, these might find themselves in exploitative schemes to provide low-cost labor for wealthier owners of NFTs [12].
+In lieu of the community-based scholarships of Axie Infinity, ICE Poker integrates elements of P2E games by design, allowing players to earn tokens just by participating in certain events, and providing them with sanctioned asset delegation.
+Wealthy owners are thus offered an official mechanism to delegate their DG wearables to willing players for whom the initial investment is perhaps prohibitively expensive.
+IV.   Methods
 For our study, we focus on the two years after the launch of Decentraland wearables on Polygon, that is, from June 2021 to May 2023.
-We excluded the short-lived first version of wearables because these were already deprecated by the launch of ICE Poker.
-In the following, we describe the overall approach to answer our RQs and then we detail how we performed the respective data collection.
-4.1 Overall approach
+For this, we undertake a data-driven approach based on public data available from blockchains and related datasets, focusing on obtaining macro-level patterns of actual user behavior related to our RQs.
+In the following, we describe the overall approach to answer our RQs and how we performed data collection.
+A.   Overall approach
 For RQ1 (wearable transfers) we first retrieved all wearables created during the period of study, marking those created by DG based on metadata from Decentraland.
-We then retrieved from the Polygon blockchain all the transfers related to all wearables, as well as their sales made in the most prominent marketplaces.
-We then analyzed the proportion between DG wearables and the rest in terms of the number of transfers as well as sales volume (both in number and USD).
-We also analyzed the share of transfers related to upgrades, which we retrieved from the respective log events on Polygon.
-For RQ2 (in-world visits to DG poker casinos), we first collected in-world tracking data of users, available only from 12 July 2022.
-Hence, we analyzed the last 10 months of the two-year period of interest, in terms of the number of accounts and time spent visiting the parcels in which DG poker casinos are located, compared to the rest of Decentraland’s map.
+We then retrieved from the Polygon blockchain the respective transfers, as well as sales and wearable upgrades from related sources.
+We then analyzed the proportion between DG wearables and the rest in terms of the number of transfers and sales volume (both in number and USD).
+For RQ2 (in-world visits to ICE poker casinos), we first collected in-world tracking data of Decentraland users, available only from 12 July 2022.
+Hence, we analyzed the last 10 months of the two-year period of interest, in terms of the number of accounts and time spent visiting the parcels in which ICE poker casinos are located, compared to the rest of Decentraland’s map.
 Our descriptive and statistical analyses mostly concern proportions, i.e., the share of wearables or land associated with DG compared with the rest of Decentraland.
-For RQ1, this is either for the whole two-year period or on a monthly basis therein.
-For RQ2, this is either for a ten-month period or on a daily basis, as we have a narrower time frame but a finer time resolution for data of in-world visits.
 To test for significant differences in proportions we used two-sided tests with significance level $\alpha=0.05$.
-When analyzing samples we used $\chi^{2}$ tests, either goodness-of-fit test or association test according to the case; when analyzing populations we used a binomial test; for both kinds of test we use relative risk (RR) to measure the effect size.
-For daily in-world visits, we also used iterative STL (i.e., seasonal and trend decomposition using LOESS) to analyze the time series, including trend and seasonal strengths (from 0 to 1).
-4.2 Data collection
+When analyzing samples we used $\chi^{2}$ tests, either goodness-of-fit test or association test according to the case; when analyzing populations we used a binomial test; for both kinds of test we use relative risk (RR) to measure effect sizes.
+B.   Data collection
 For wearable metadata (e.g., collection, category, creator), we used several of the openly available Decentraland’s query endpoints on TheGraph, a service for querying blockchain networks, and whose definitions are open source.
-[^1] https://github.com/decentraland For transfers (including mints), we collected the respective transactions from Polygon via the blockchain web service Alchemy.
-For wearable sales, we collected data from two different sources: 1) Decentraland’s Marketplace for internal sales, using the aforementioned TheGraph query endpoints; and 2) OpenSea, the largest overall NFT marketplace and main means for external sales, as well as the initial default marketplace for DG wearables.
+[^1] https://github.com/decentraland For transfers, we collected the respective transactions from Polygon via the blockchain web service Alchemy.
+For sales, we collected data from two different sources: 1) Decentraland’s Marketplace for internal sales, using the aforementioned TheGraph query endpoints; and 2) OpenSea, the largest overall NFT marketplace and main means for external sales, as well as the initial default marketplace for DG wearables.
 We also retrieved from Yahoo!
-Finance the price of the few different cryptocurrencies used for the collected wearable sales during the two-year period of study.
-To normalize the price of sales, we first converted it to USD, then we adjusted it for inflation on a monthly basis as of December 2023, based on the Consumer Price Index (CPI) by the Bureau of Labor Statistics of the United States.
-Price changes of these cryptocurrencies also serve to gauge the interest on the studied platforms.
-Indeed, as can be seen in Figure 4, the ICE token was released around Meta’s rebranding in November 2021, after the latter provoked a sudden price increase in Decentraland’s MANA (fourfold in just a couple of days).
-This was a considerable surge, even taking into account the already upward trend of MATIC in the wider Polygon ecosystem.
-However, the collapse of several interconnected cryptocurrencies the following May marked the onset of the 2022 cryptowinter OECD (2022), which affected a large share of blockchain projects and lasted until mid 2023.
-Figure 4: Cryptocurrency price in USD during the two-year period of study for: ICE by Decentral Games, MANA by Decentraland, and MATIC by Polygon.
-For data on DG wearable upgrades, we again used Alchemy to retrieve the log events for the period of study of DG’s smart contract on Polygon [^2] Address 0xC9a67eD1472A76d064C826B54c144Ca00DAE4015 which manages the respective on-chain transactions, called “iceRegistrant”.
+Finance the price of the few cryptocurrencies used in the collected wearable sales, which we adjusted for inflation.
+Price changes of these cryptocurrencies also serve to gauge interest on the studied platforms.
+As can be seen in Figure IV, the ICE token was released around the rebranding of Facebook into Meta in November 2021, with the latter provoking a sudden price increase in Decentraland’s MANA (fourfold in just a couple of days).
+This was a considerable surge, even taking into account the already upward trend of MATIC, Polygon’s cryptocurrency.
+However, the collapse of several interconnected cryptocurrencies the following May marked the onset of the 2022 cryptowinter [42], which affected a large share of blockchain projects and lasted until mid 2023.
+Figure IV: Cryptocurrency price in USD during the two-year period of study for: ICE by Decentral Games, MANA by Decentraland, and MATIC by Polygon.
+For data on wearable upgrades, we again used Alchemy to retrieve the log events of the DG’s smart contract on Polygon [^2] Address 0xC9a67eD1472A76d064C826B54c144Ca00DAE4015 that managed such transactions.
 We then used the aforementioned wearable metadata to derive the ranking level and rarity limit of the upgraded tokens.
-For in-world user tracking, we utilized the corresponding data archives from Atlas Data Corp, a blockchain analytics company that obtained a grant  [^3] https://decentraland.org/governance/project/?id=7422a99e-8a25-4625-9b30-ab688de5dade from Decentraland’s DAO to finance a system to keep record of user sessions.
+For in-world user tracking, we utilized the corresponding data archives from Atlas Data Corp, a blockchain analytics company that obtained a grant  [^3] https://decentraland.org/governance/project/?id=7422a99e-8a25-4625-9b30-ab688de5dade from Decentraland’s DAO to keep record of user sessions.
 These data are log events (created every circa 20 seconds) of the current location of every connected user on the several servers that manage the 3D rendering of in-world scenes and avatars on Decentraland.
 Given that this system started operating in mid July 2022, we narrowed our analysis of in-world visits to only 10 months, from August 2022 to May 2023.
-Moreover, we take into consideration only users with an Ethereum-compatible address, that is, we discard the small minority of guest users not associated with an account.
+Moreover, we take into consideration only users with an Ethereum-compatible address, that is, we discard the small minority of guests not associated with an account.
 We then transformed the more than 677 million log events for this time frame into sessions with a start and end for each land parcel and account address.
 Finally, we segmented the respective sessions’ duration on a daily basis, adding 10 seconds to each to compensate for possible shortenings due to the 20-second polling rate of the log events.
-5 Analyses and Results
-5.1 RQ1: Overall wearable transfers
+V.   Analyses and Results
+A.   RQ1: Overall wearable transfers
 From June 2021 to May 2023, a total of 4,629 collections with 8,450 wearable items (i.e., designs) were published by 1,517 distinct creators.
-Of these, we identified and manually verified two accounts [^4] Addresses 0x17a253c2ac0d5ba92cadbbf665e3390c9913dc5d and 0x00e5d44f6a296c10f159486f842838bd68f13e32 that created all of the DG wearables, based on marketplace offerings, item names, and user profiles on Decentraland.
+Of these, we identified and manually verified two accounts [^4] Addresses 0x17a253c2ac0d5ba92cadbbf665e3390c9913dc5d and 0x00e5d44f6a296c10f159486f842838bd68f13e32 that created all of DG wearables, based on marketplace offerings, item names, and user profiles on Decentraland.
 Both DG accounts (0.1% of creators) published 117 collections with 723 items (8.6% of the total available items).
@@
 […]
-In our analysis on a monthly basis, we noticed great variation in the share of DG wearables over time, but interestingly, the variations are different for transfers and sales.
-On one hand, and as illustrated in Figure 5, the total volume of transfers reached its highest values in March and April of 2022, which coincides with the first edition of the Metaverse Fashion Week at the end of March.
+In our analysis on a monthly basis, we noticed great variation in the share of DG wearables over time, but in different ways for transfers and sales.
+On one hand, and as illustrated in Figure V, the total volume of transfers reached its highest values in March and April of 2022, which coincides with the first edition of the Metaverse Fashion Week at the end of March.
 In those two months the share of DG wearables was only 3.3% and 4.5%, respectively.
@@
 This DG share increase is mostly related to a few wearable items with a common rarity limit and which give access only to ICE Poker Arcade, whose app was released precisely at the end of September.
-On the other hand, for the monthly sales volume in USD (illustrated in Figure 6), the highest monthly value for both total volume and share of DG wearables (62.9%) is November 2021, the month in which the first DG collections focused on ICE Poker were released for minting.
-Moreover, the timing also coincides with the wake of Meta’s rebranding, which provoked a surge and sustained high price in both ICE and MANA for that month, as shown in Figure 4.
-Figure 5: Monthly number of transfers of Decentraland wearables on the Polygon blockchain and respective share associated with Decentral Games (DG).
-The upward trend in the bottom panel is related to wearables that grant access to the ICE Poker Arcade app, released in September 2022.
-Figure 6: Monthly sales volume in USD of Decentraland wearables —made via the marketplaces of Decentraland and OpenSea— and the respective share associated with Decentral Games (DG).
-The bottom panel’s spike in November 2021 is related to the release of ICE-themed DG wearables and the high valuation of ICE in the wake of Meta’s rebranding.
+On the other hand, for the monthly sales volume in USD (illustrated in Figure VI), the highest monthly value for both total volume and share of DG wearables (62.9%) is November 2021, the month in which the first DG collections focused on ICE Poker were released for minting.
+Moreover, the timing also coincides with the wake of Meta’s rebranding, which provoked a surge and sustained high price in both ICE and MANA for that month, as shown in Figure IV.
 For the two-year period there were 58,791 mints of wearables with an ICE ranking level, of which 42,390 (72.1%) were via upgrades.
@@
 In other words, 90% of owners preferred to burn a lower-ranking token and mint a higher-ranking token via an upgrade, rather than buy (or wait to win) one without sacrificing the former.
-5.2 RQ2: In-world visits of poker casinos
+Figure V: Monthly number of transfers of Decentraland wearables on the Polygon blockchain and the respective Decentral Games (DG) share.
+Figure VI: Monthly sales volume in USD of Decentraland wearables —made via the marketplaces of Decentraland and OpenSea— and the respective share associated with Decentral Games (DG).
+B.   RQ2: In-world visits of poker casinos
 […]
-When analyzed on a daily basis, visits to the ICE Poker casinos are subject to relatively small fluctuations —compared to the rest of the world map— both in terms of time spent (see Figurge 7) and number of distinct visitor accounts (see Figure 8).
-For visits to ICE Poker casinos, we have a moderate but steady downward trend for both time spent and visitors, with a trend strength (based on STL decomposition) of 0.98 and 0.99, respectively.
-In the rest of the world map, the trend for time spent is slightly downward and with a strength of 0.83 for time spent, while the trend is visibly downward with a strength 0.91 for visitors.
+When analyzed on a daily basis, visits to the ICE Poker casinos are subject to relatively small fluctuations —compared to the rest of the world map— both in terms of time spent (see Figure VII) and number of distinct visitor accounts (see Figure VIII).
+For visits to ICE Poker casinos, we have a moderate but steady downward trend for both time spent and visitors.
+In the rest of the world map, the trend is slightly downward for time spent, while the trend is visibly downward for visitors.
 In other words, over time there were less distinct accounts visiting Decentraland, but these collectively spent more time on it.
 However, upon further inspection this increase in time spent per visitor is mostly due to a few accounts with a much higher number of hours than average.
-Indeed, in terms of share, time spent on ICE Poker casinos reached its highest levels at the beginning of September 2022, up to 53%, but it decreased to circa 20% in May 2023 due the the aforementioned outlier users in the rest of the world map.
+Indeed, in terms of share, time spent on ICE Poker casinos reached its highest levels at the beginning of September 2022, up to 53%, but it decreased to circa 20% in May 2023 due to the aforementioned outlier users in the rest of the world map.
 Nevertheless, for visitors, the share at the beginning of the ten-month period was of circa 32%, and at the end it increased to circa 41%.
-Curiously, the number of visitors of ICE Poker casinos had a modest weekly seasonality strength of 0.66, with the number of visitors slightly decreasing on weekends; for the other time series the seasonality strength was less than 0.5.
-Figure 7: Stacked area chart of daily time spent (in hours) visiting Decentraland.
+Figure VII: Daily time spent (in hours) visiting Decentraland.
 ICE Poker casinos had a remarkable share, with a daily median of 20%.
 The highest overall spike was reached during the Metaverse Music Festival (MVMF) in mid-November.
-Figure 8: Stacked area chart of daily distinct accounts that visited Decentraland.
+Figure VIII: Daily distinct accounts that visited Decentraland.
 ICE Poker casinos had a striking share, with a daily median of 33%.
 The Metaverse Music Festival (MVMF) in mid-November had the highest spike overall, while the Metaverse Fashion Week (MVFW) at the end of March had the only spike during the last four months.
-These trend and seasonality strength differences and shares by land category are in part explained by their dissimilar spikes and surges.
 On one hand, visits to the ICE Poker casinos had no spikes nor dips.
@@
 In particular, the Decentraland Metaverse Music Festival, held on 10–13 November 2022, was the event that provoked the most visits (both in time and accounts).
-The festival was free to attend, and its line up involved famous artists such as Björk, Ozzy Osbourne, Dillon Francis, and Soulja Boy.
 Incidentally, the Metaverse Fashion Week of 2023 also provoked a spike in visitors at the end of March, but not in time spent and only during a period of four months in which the number of distinct visitors had already remarkably declined.
-6 Discussion
-Based on our results, we can respond affirmatively to both RQ1 and RQ2: wearable transfers on Decentraland are indeed disproportionately related to those created by DG, and the ICE Poker casinos do receive significantly (and remarkably) more visits than other in-world locations.
-In the following, we first discuss our main findings concerning the influence of DG and ICE Poker on Decentraland, we comment on the initiatives and events affecting such influence over time, we reflect upon the ambiguous legal standing of ICE Poker, we explore the social implications of our findings, and, finally, we describe the main limitations of our study.
-6.1 Main findings
+VI.   Discussion
+Based on our results, we can respond affirmatively to both RQ1 and RQ2: wearable transfers on Decentraland are indeed disproportionately related to those created by DG, and ICE Poker casinos do receive significantly (and remarkably) more visits than other in-world locations.
+In the following, we discuss our main findings, we reflect upon the ambiguous legal standing of ICE Poker, we explore the social implications of our findings, and finally, we describe the main limitations of our study.
+A.   Main findings
 Our findings confirm that DG and ICE Poker have great influence on Decentraland wearable and visit dynamics.
-Just two DG accounts out of 1,517 wearable creators published 8.6% of the total available items, with these being twice more likely to be minted (i.e., involved in a creation transfer) based on the maximum supply limit compared to the rest, particularly at the end of the period of study.
+Just two DG accounts out of 1,517 wearable creators published 8.6% of the total available items, with these being twice more likely to be minted based on the maximum supply limit compared to the rest, particularly at the end of the period of study.
 DG wearables also had a monetary sales volume an order of magnitude higher despite the considerably fewer sales.
@@
 The initiative, however, did not meet expectations.
-Based on a qualitative study on the 2023 edition Miikkulainen-Gilbert (2023), most user complaints were related to the on-boarding complexity of setting up an Ethereum account, a lackluster in-world user experience, the context switch between in-world and traditional web-based interaction for many events, the technical limitations of wearable 3D modeling, and the need of paid approval even for wearables designed for oneself.
-In our analysis of visits we indeed detected a spike of distinct visitors during the 2023 edition at the end of March (see Figure 8), but there was no visible effect on the time spent on the platform (see Figure 7), and the number of visitors continued its downward trend afterward.
+Based on a qualitative study on the 2023 edition [38], most user complaints were related to on-boarding complexity, a lackluster in-world user experience, the context switch between in-world and traditional web-based interaction, and the need of paid approval even for wearables designed for oneself.
+In our analysis of visits we indeed detected a spike of distinct visitors during the 2023 edition at the end of March (see Figure VIII), but there was no visible effect on the time spent on the platform (see Figure VII), and the number of visitors continued its downward trend afterward.
 The fact that in 2024 there was no Metaverse Fashion Week confirms the failure of the initiative to keep up the momentum going.
-On the other hand, the Metaverse Music Festival attracted much more visitors that spent considerable time in Decentraland (see Figures 7 and 8), but nevertheless visits did not visibly increase in the following weeks.
-Unlike in-world fashion events, music concerts have been present for many years in social virtual worlds and MMO games, and during the Covid-19 pandemic they were seen as an attractive alternative for the struggling live music industry Çeli̇k (2023).
-It remains to be seen if this Decentraland initiative will continue to gather interest in future editions.
+On the other hand, the Metaverse Music Festival attracted much more visitors that spent considerable time in Decentraland (see Figures VII and VIII), but nevertheless visits did not visibly increase in the following weeks.
+It remains to be seen if this initiative will continue to gather interest.
 As for DG, the main initiative after the release of ICE Poker wearables was the introduction of Tournament mode, which changed the rules to win prizes and claim ICE tokens.
-These changes were deemed necessary due to the quick pace with which delegates, who mostly did not own other kinds of DG assets (e.g., wearables or its utility token), withdrew ICE to convert it to other cryptocurrency or fiat money, which accelerated its inflation.
-Based on a financial report report by an external firm made in collaboration with DG, the aforementioned changes achieved their goal by reducing the inflation of ICE by circa 58% Sun (2023).
-In order to not alienate core delegates with such changes, DG gifted wearables to the 3,000 most active ones, which seems to explain the surge in visitors and time spent in ICE Poker casinos at the end of August 2022 (see Figures 7 and 8).
-Another important DG initiative was the launch of its Arcade mode and related web and mobile apps in September 2022, which allow users to play and earn non-monetary prizes without the need of visiting Decentraland; possession of DG wearables is still necessary, however.
-This explains the increase in the share of transfers of DG wearables at the end of the period of study (see Figure 5) without the corresponding increase in visits share to ICE Poker casinos.
+These changes were deemed necessary due to the quick pace with which delegates, who mostly did not own other kinds of DG assets (e.g., wearables or its utility token), withdrew ICE to convert it to other cryptocurrency or fiat money, accelerating its inflation.
+In order to not alienate core delegates with these changes, DG gifted wearables to the 3,000 most active ones, which seems to explain the surge in visitors and time spent in ICE Poker casinos at the end of August 2022 (see Figures VII and VIII).
+Another important DG initiative was the launch of its Arcade mode and related web and mobile apps in September 2022, which allow players to earn non-monetary prizes without the need of visiting Decentraland; possession of DG wearables is still necessary, however.
+This may explain the increase in the share of transfers of DG wearables at the end of the period of study (see Figure V) without the corresponding increase in visits share to ICE Poker casinos.
 Concerning external market forces, the ICE token and the initial ICE Poker wearables were by happenstance released a few days shy of Meta’s rebranding.
-This image change of such a behemoth corporation is considered to be the main factor behind the surge in popularity of the term Metaverse and the hype thereof Dolata & Schwabe (2023).
-Consequently, the valuation of DG’s newly introduced cryptocurrency and wearables reached prominent heights, along with Decentraland’s own cryptocurrency, as can be seen in Figures 4 and 6.
-Nevertheless, the initial hype regarding NFTs has subsided (and to a lesser extent also that of the Metaverse), with attention having shifted to generative artificial intelligence, which came to the spotlight with the release of ChatGPT Echauri (2024).
-In addition, platforms based on tokenomics must continually adapt to the hectic blockchain community, volatile cryptocurrencies, and stricter regulations OECD (2022).
-Indeed, at the moment of writing (early 2024), DG is transitioning into a single utility and cryptocurrency token called BAG, to simplify its tokenomics and counter some of the remaining inflationary issues of ICE Decentral Games (2024).
-DG also plans to revamp its games, sunsetting the current system of upgrades and delegations in favor of improving the utility of owning a DG wearable, and it is also consolidating its operating license, which is currently on questionable legal standing.
-6.2 On the legality of ICE Poker
-After a few years of uncertainty regarding its legal status, online gambling is now banned or heavily regulated in most jurisdictions Zborowska et al.
-(2012).
-Similarly, for many years and due to their novelty, blockchain cryptocurrencies remained in a regulatory limbo in most jurisdictions Yeoh (2017).
-However, after many instances of scams and fraud —some very large and highly mediatized, as in the case of the exchange FTX— in the last few years there has been a push to outright ban or closely regulate cryptocurrencies Trautman et al.
-(2022).
-ICE Poker is a combination of both gambling and cryptocurrencies, hence, this begs the simple question on whether it is a legal operation or not.
+This image change of such a behemoth corporation is considered to be the main factor behind the surge in popularity of the term Metaverse and the hype thereof [13].
+Consequently, the valuation of DG’s newly introduced cryptocurrency and wearables reached prominent heights, along with Decentraland’s own cryptocurrency, as can be seen in Figures IV and VI.
+Nevertheless, the initial hype regarding NFTs has since subsided (also that of the Metaverse), with attention having shifted to generative artificial intelligence, which came to the spotlight with the release of ChatGPT [15].
+B.   On the legality of ICE Poker
+After a few years of uncertainty regarding its legal status, online gambling is now banned or heavily regulated in most jurisdictions [57].
+Similarly, for many years and due to their novelty, cryptoassets remained in a regulatory limbo in most jurisdictions, but after many instances of scams and fraud —as in the case of the exchange FTX— in the last few years there has been a push to outright ban or closely regulate cryptocurrencies [50].
+ICE Poker is a combination of both gambling and cryptocurrencies.
+Hence, this begs the simple question on whether this kind of operation is legal or not.
 The answer is, however, murky.
-Firstly, on the project’s front page it is stated that DG is operated by the company Kidderminster Cheltenham, whose official address is in Costa Rica.
-This is most likely due to the country’s laissez-faire approach for both online gambling and cryptocurrencies  Kerr (2018), its tax haven tendencies Baron (2019), and the fact that its operating licenses are often used to offer an online gambling service in foreign jurisdictions in which the licensing process is more rigid or non-existing Fischer (2021).
-Moreover, in the official project documentation it is stated that the company behind DG is called Web4 LTD, which is established in the British Virgin Islands, the place of jurisdiction regarding the terms and conditions of the project, and a British overseas territory known for being a tax haven and having links to dubious online gambling operations Macegoniuk (2021).
-Curiously, in the project’s website there is also a legal opinion regarding ICE Poker by a consultant from the Singaporean firm Jacque Law Zhirong (2022), in which it is stated that the game “would not constitute ‘gambling’, ‘betting’, ‘lottery’ under the RGA [Singaporean Remote Gambling Act of 2014]” because “gambling laws are not intended to cover games whereby there is no staking or wagering of money or money’s worth”.
-Indeed, possession of a DG wearable is necessary to earn ICE, but the wearable has independent value outside of ICE Poker and its use is not exclusive to DG.
-In addition, users play with chips, which according to the opinion Zhirong (2022), “are non-transferable tokens that hold no value or functionality other than to keep score when players play ICE Poker”.
-Based on the number of chips and wearables they possess, players periodically receive ICE, which can be converted to fiat money, but on an external exchange such as Binance.
-This convoluted mechanism is similar to the one used offline by pachinko parlours in Japan to circumvent gambling laws Brooks et al.
-(2008): pachinko players win small steel balls, which can be exchanged for a receipt with a nominal value, which in turn is used to claim a low-valued physical item (e.g., a specific pack of washing powder), and finally, the player can exchange this item for a lesser nominal value in cash outside the parlour on a nearby “independent office”.
-Of course, the above legal opinion applies only to Singapore.
-In the case of other jurisdictions DG is more cautious, claiming that the onus is on players.
-In fact, to participate in its virtual casinos, players must guarantee they are not residents of one of the non-supported regions listed on the project’s website.
-At the moment of writing (early 2024), there were 36 non-supported regions, including high-income countries such as the United States, the United Kingdom, France, Germany, and Australia.
-ICE Poker Arcade might nonetheless be supported in these regions, given that users play to win physical or digital prizes, not money or cryptocurrency.
-To address the issue in a more solid manner, recently DG acquired a 90% stake in Virtue Gaming Ltd, the entity that holds a Malta license to enable real-money gambling in Virtue Poker, the first licensed cryptocurrency-based online poker.
-DG intends to expand this coveted license to its virtual casinos and other games, which would remove doubts regarding their current legal status in many jurisdictions.
-6.3 Social implications
-Unlike most previous social virtual worlds and social media in general, platform transparency and openness are stated to be paramount in Decentraland.
+Firstly, on the project’s front page it was stated that DG’s parent company operated in Costa Rica.
+This is most likely due to the country’s laissez-faire approach for both online gambling and cryptocurrencies  [34], and the fact that its operating licenses are often used for online gambling services in foreign jurisdictions in which licensing is more rigid or non-existing [19].
+Moreover, in the official project documentation it is stated that the company developing DG, Web4 LTD, is established in the British Virgin Islands, a British overseas territory known for being a tax haven and having links to dubious online gambling operations [37].
+Curiously, the terms and conditions of ICE Poker state that is not covered by gambling laws in many jurisdictions due to the lack of (fiat) money exchange therein.
+Based on the number of chips and wearables they possess, players periodically receive ICE, which can ultimately be converted to fiat money, but on an external exchange such as Binance.
+Yet, players must guarantee they are not residents of one of the dozens of non-supported regions listed on the project’s website due to their strict gambling and cryptocurrency regulations.
+C.   Social implications
+Platform transparency and openness are stated to be paramount in Decentraland.
 The project strives toward these qualities by means of its use of public records on the blockchain, user-oriented ownership of in-world assets, public and open-source code repositories, governance via a DAO, and initiatives to overtly monitor the ecosystem —e.g., by financing the collection of in-world tracking data we used to analyze visits.
-These characteristics, common to many blockchain-based projects, also offer unprecedented opportunities for both industry and academia to develop and understand novel sociotechnical systems, in the form of virtual reality spaces but not only Cannavo & Lamberti (2020).
-Nevertheless, the current levels of transparency and openness in Decentraland do not preclude the abuse of the platform’s mechanisms.
+These characteristics, common to many blockchain-based projects, also offer unprecedented opportunities for both industry and academia to develop and understand novel sociotechnical systems in the form of virtual reality spaces, but not only [9].
+Nevertheless, the current levels of transparency and openness in Decentraland do not preclude the abuse of the platform.
 For instance, in June 2022 DG submitted a proposal to Decentraland’s DAO for a grant, [^5] https://decentraland.org/governance/proposal/?id=99c66b90-e2d2-11ec-9000-175d8dd584b8 asking the equivalent to one million USD in MANA to increase its liquidity pool for ICE rewards.
@@
 This is an illustrative example of how a decentralized governance is not necessarily a neutral one.
-Rather, in its current form Decentraland’s DAO gives way to automated situations reminiscent of the principal–agent problem Goldberg & Schär (2023), because “what is best for a dominant individual is not necessarily a good choice for the virtual world as a whole as well as the majority of its stakeholders”.
-Other examples of potential platform abuse are the aforementioned drastic reward changes introduced with Tournament mode and the current revamping of ICE Poker itself.
+Rather, in its current form Decentraland’s DAO gives way to automated situations reminiscent of the principal–agent problem [22], because “what is best for a dominant individual is not necessarily a good choice for the virtual world as a whole as well as the majority of its stakeholders”.
+Other examples of potential platform abuse are the aforementioned drastic reward changes introduced with Tournament mode and the revamping of ICE Poker itself.
 These highlight that despite the assurances of public transactions and smart contracts, most of the functionality of the poker system —and the virtual world itself— lives outside the blockchain, and is thus subject to arbitrary changes by its controlling entity.
 In addition, a third-party actor was able to exploit the affordances provided by the virtual world, which lead to deviations in user behavior unforeseen by the platform’s creators.
-The decision by DG to use Decentraland wearables as access tokens to play in its online casino games has significantly influenced the overall acquisition trends of wearables at the platform level, tilting the motivating factors for acquiring wearables from intrinsic (e.g, the aesthetic enjoyment of the garment) to extrinsic (e.g., the earning utility of the token).
-Hence, functional wearables might be appreciated solely for their utility, whereas cosmetic ones might be deemed less valuable due to their absence of utility, which weakens Decentraland’s claim to be a creative social space.
-Besides, based on our results in-world visits seem be greatly influenced by the earning potential within ICE Poker casinos, instead of the ludic enjoyment that these and other world areas offer.
-However, an overabundance of extrinsic motivation can often undermine an experience by eroding the spontaneous intrinsic motivation, especially for games Delic & Delfabbro (2024); Hong et al.
-(2023).
-At the same time, in online poker the earning prospects are partially integrated into the gameplay, thus becoming part of the intrinsic motivation to play, but with the possibility of leading to problem gambling without the proper containment mechanisms Zaman et al.
-(2014).
-ICE Poker sits at the intersection of gamblers and cryptoasset traders, with potentially troubling consequences as both profiles might exhibit similar self-destructive behavior, such as engaging in risk-taking and thrill-seeking activities, and a susceptibility to addiction Johnson et al.
-(2023); Delfabbro et al.
-(2021b).
-We should not forget, however, that in Ice Poker, as with other play-to-earn games, a considerable amount of grunt play is done by low-income people who do not own assets and for whom game enjoyment is secondary.
-It is a new form of digital exploitation —where the lion’s share goes to the owner of the NFT who lends it to the less fortunate— and which could lead to increased social disparity and potentially exacerbate economic inequalities Zaucha (2024).
-For all the above reasons, the study of the emerging phenomena in novel social platforms such as Decentraland and games such as Ice Poker is increasingly important.
-A better understanding of their underlying mechanisms and dynamics will allow, for instance, polycymakers to be better informed and prepared to enact the necessary changes to tackle these and other related issues that might emerge in the next few years.
-6.4 Limitations
+The decision by DG to use Decentraland wearables as access tokens to play in its online casino games has significantly influenced the overall acquisition trends of wearables at the platform level, perhaps tilting the motivating factors for acquiring wearables from intrinsic (e.g, the aesthetic enjoyment of the garment) to extrinsic (e.g., the earning utility of the token).
+Hence, functional wearables might be appreciated solely for their utility, whereas merely cosmetic ones might be deemed less valuable due to their absence of utility, which weakens Decentraland’s claim to be a social space primarily focused on creativity and self-expression.
+In fact, an overabundance of extrinsic motivation can often undermine an experience by eroding the spontaneous intrinsic motivation, especially for games [29].
+Based on our results, in-world visits are also greatly related to DG casinos.
+As with wearable transfers, this phenomenon is likely influenced by the earning potential of ICE Poker, in addition to the ludic enjoyment alone that casinos and other world areas offer.
+Moreover, ICE Poker combines gambling and cryptoasset trading, with potentially troubling consequences to players due to self-destructive behavior, such as engaging in risk-taking and thrill-seeking activities, and a susceptibility to addiction [31].
+Lastly, professional gamblers in ICE Poker and ICE Arcade could potentially replicate exploitative practices observed in other Play-to-Earn (P2E) games.
+They might use scholarships and delegation as a form of digital feudalism, where the majority of the profits are funneled to a small group of individuals or entities.
+This could lead to increased social disparity and potentially exacerbate economic inequalities [56], particularly if low-income individuals are involved, similar to the situation observed in Axie Infinity [10].
+The study of the emerging phenomena in novel platforms such as Decentraland and games such as Ice Poker is increasingly important.
+A better understanding of their underlying mechanisms and dynamics will allow stakeholders to be better informed and prepared to enact the necessary changes to tackle these and other related issues that might emerge in the next few years.
+In particular, we invite policymakers to take a critical look into the growing phenomenon of gambling and gambling-like activities in social virtual worlds based on tokenomics.
+Given the risk-taking behavior associated with both cryptocurrencies and gambling, together with their novel use in immersive virtual worlds, consumer protection should be a priority, especially for vulnerable individuals, such as those who are both low-income and younger.
+To this end, policymakers could follow an approach similar to the recent tackling of abuse of loot boxes in video games [54], by expanding the legal definition of gambling and adapting existing regulation, such as enforcing the disclosure of probabilities regarding risks and rewards, and applying age rating and warning labels.
+In view of the above legal and economic uncertainties, we also suggest developers of social virtual worlds to be wary of heavily relying —directly or indirectly— on such gambling-like activities to keep users engaged.
+Moreover, particular care should be taken in designing and implementing the decentralized governance and affordances of the platform, as to not allow a single third party to gain an overtly dominant role.
+Indeed, Decentraland’s reliance on ICE Poker represents a predicament for the platform, which already suffers from a volatile ecosystem and diminished interest, raising doubts on its long-term stability.
+In this regard, at the end of 2024 the project announced a new version of the platform, Decentraland 2.0, with the stated goals of strengthening social interactions, boosting user engagement, improving discoverability, and increased immersion and fun.
+At the moment of writing, it remains unclear if this new version will succeed in reaching its goals and if it will overcome the dominance of in-world gambling activities.
+D.   Limitations
 In spite of our careful approach, several limiting factors could influence the interpretation and applicability of our results.
@@
 Moreover, these might even change over time.
-For instance, we noticed a sales drop that lasted several days on OpenSea in May 2022, which upon further inspection resulted from a migration to their current protocol called SeaPort.
 Nonetheless, the marketplaces we selected should cover the vast majority of wearable sales, as Decentraland’s Marketplace is the official and default venue, and OpenSea is the largest NFT marketplace and was the initial default venue to sell wearables by DG, which now has its own marketplace.
@@
 We did not use the term active users in our analysis because we only considered the presence of a given account address in the log events, i.e., we did not filter out potential bots or users that could be deemed inactive.
-In line with other kinds of social media, bots are also a plague in social virtual worlds, with automated avatars having a noticeable impact on these Varvello & Voelker (2010).
+In line with other kinds of social media, bots are also a plague in social virtual worlds, with automated avatars having a noticeable impact on these [52].
 In addition, a user might be authentic (i.e, not a bot) but be away from keyboard while their avatar is still present in the virtual world.
 Determining an avatar who represents an “authentic active user”, however, depends on an arbitrary set of characteristics and mechanisms.
-Indeed, calculating the number of daily active users on Decentraland has sparked much public debate, as different definitions and mechanisms might result in vastly different numbers Thompson et al.
-(2022).
-Our approach, on the other hand, is relatively simplistic, but at the same time it better responds RQ2, as in any case these avatars —active or inactive, human or bot— occupy space and resources on Decentraland’s virtual world.
-7 Conclusions
+Our approach, on the other hand, is relatively simplistic, but at the same time it better answers RQ2, as in any case these avatars —active or inactive, human or bot— occupy space and resources on Decentraland’s virtual world.
+Finally, our focus on quantitative methods and the lack of qualitative methods, such as user interviews and questionnaires, could be a limitation for investigating the motivations behind users’ actions.
+We believe that using a quantitative approach based on public data to identify global patterns of actual user behavior is appropriate to answer our study’s RQs.
+Nevertheless, future studies could also adopt a more comprehensive and multifaceted approach using mixed methods.
+For instance, by corroborating data analyses such as ours with surveys to obtain data about the users’ self-perceived behavior, or even with ethnographic analyses [5, 8] to delve into the details of users’ behaviors in social virtual worlds.
+VII.   Conclusions
 The concept of the Metaverse has in recent years sparked the collective imagination on novel and creative ways in which we could interact within immersive social virtual worlds and represent ourselves via our avatars.
@@
 This predominance of a single third-party online gambling initiative —which has transformed itself several times due to technical, economic, and legal uncertainties— raises doubts on if and how Decentraland’s current bet on tokenomics to realize its vision of the Metaverse will ultimately pay off.
+Funding
+This research received no specific grant from any funding agency in the public or private sectors.
+Authors’ Contributions
+All authors have participated in drafting the manuscript. All authors read and approved the final version of the manuscript.
+Conflict of Interest
+The authors declare no conflict of interest.
+Data Availability
+The data supporting the findings of this study are available on FigShare at: https://doi.org/10.6084/m9.figshare.28596923.
+Ethical Statement
+Ethical approval was not required for the present study due to its use of anonymous and public data.
+In particular, the vast majority of our data was obtained from blockchain platforms with permissive software and data licenses.
+For these data, there is not personally identifiable information of the analyzed users, as these are anonymously identified via a unique string of hexadecimal characters that was cryptographically generated.
+Declaration of AI Usage
+No AI tools were used in the creation of this manuscript.